### v3

# Random Sperner lemma and random Brouwer fixed point theorem

Abstract
In this paper, we systematically develop the theory of $L^{0}$-simplicial subdivisions of $L^{0}$-simplexes by grasping the inherent connection between $L^{0}$-simplexes and usual simplexes. We establish a key representation theorem of a proper $L^{0}$-labeling function by usual proper labeling functions. Based on the representation theorem, we can establish a general random Sperner lemma. Next, as an application of the random Sperner lemma, we give a complete proof of the known random Brouwer fixed theorem in random Euclidean spaces. Besides, all the works currently available and closely related to the randomization of the classical Brouwer fixed point theorem are unified.
keywords:
Random Sperner Lemma, Random Brouwer fixed point theorem, Random Euclidean space, $L^{0}$-module, $L^{0}$-simplex
pacs:
MSC Classification]46H25, 47H10, 46A50, 60H25
1 Introduction
Throughout, $\mathbb{R}^{d}$ denotes the $d$-dimensional Euclidean space. The well known Brouwer fixed point theorem [3] states that every continuous mapping from a bounded closed convex subset of $\mathbb{R}^{d}$ to itself has a fixed point. Brouwer fixed point theorem has rich applications in analysis, geometry, differential equations and mathematical economics [39, 8, 14, 40], in particular, it is also attractive that Brouwer fixed point theorem is equivalent to many famous fundamental results [38, 29, 30, 41], for example, the Knaster-Kuratowski-Mazurkiewicz (briefly, $KKM$) mapping principle [33], the classical Borsuk theorem [2] and the Gustafson-Schmitt-Cid-Mawhin’s theorem on the existence of periodic solutions of ordinary differential equations [8, Theorem 1.1]. Due to its fundamental importance, Brouwer fixed point theorem has been generalized in various ways [32, 46, 50, 7, 13] and has had a major impact on the development of nonlinear analysis and fixed point theory [1, 9, 18].
With the aim of developing random functional analysis [23, 26], dynamic Nash equilibrium theory [11, 26] and stochastic differential equations [42, 43, 44, 45], there are strong desires to extend Brouwer fixed point theorem to the setting of a random Euclidean space $L^{0}(\mathcal{F},\mathbb{R}^{d})$, where $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is the linear space of equivalence classes of random vectors from a probability space $(\Omega,\mathcal{F},P)$ to the $d$-dimensional Euclidean space $\mathbb{R}^{d}$. Since $L^{0}(\mathcal{F},\mathbb{R}):=L^{0}(\mathcal{F},\mathbb{R}^{1})$ is an algebra over the real number field $\mathbb{R}$, $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is a free module of rank $d$ over the algebra $L^{0}(\mathcal{F},\mathbb{R})$, and thus the notions of an $L^{0}$-convex set and a $\sigma$-stable set can be introduced and developed. More importantly, $L^{0}(\mathcal{F},\mathbb{R}^{d})$ can be regarded as a special complete random normed module. The notion of a random normed module was introduced independently by Guo in connection with the idea of randomizing space theory [19, 21, 22, 23, 25, 26] and by Gigli in connection with nonsmooth differential geometry on metric measure spaces [16] (see also [4, 5, 6, 17, 36, 37] for the related advances). When regarded as a random normed module, $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is often endowed with two kinds of useful topologies: one is the $(\varepsilon,\lambda)$-topology, denoted by $\mathcal{T}_{\varepsilon,\lambda}$, which coincides with the usual topology of convergence in probability, is a typical locally nonconvex linear topology and makes $(L^{0}(\mathcal{F},\mathbb{R}^{d}),\mathcal{T}_{\varepsilon,\lambda})$ a topological module over the topological algebra $(L^{0}(\mathcal{F},\mathbb{R}),\mathcal{T}_{\varepsilon,\lambda})$; the other is the locally $L^{0}$-convex topology, denoted by $\mathcal{T}_{c}$, which was introduced in [15] and coincides with the topology generated by random open balls and makes $(L^{0}(\mathcal{F},\mathbb{R}^{d}),\mathcal{T}_{c})$ a topological module over the topological ring $(L^{0}(\mathcal{F},\mathbb{R}),\mathcal{T}_{c})$, see [23] for the relevant connections between the two kinds of topologies. Thus, according to different choices of $\mathcal{T}_{\varepsilon,\lambda}$ and $\mathcal{T}_{c}$, random Brouwer fixed point theorem has the corresponding two versions (or forms): the first version is the formulation of random Brouwer fixed point theorem under $\mathcal{T}_{c}$, which was first given in [26] and is stated as Theorem 2.1 in this paper; the second version is the formulation of random Brouwer fixed point theorem under $\mathcal{T}_{\varepsilon,\lambda}$, which was first given in [11, 28] and is stated as Proposition 2.3 in this paper. It follows immediately from Lemma 4.3 of [26] that Proposition 2.3, namely the $\mathcal{T}_{\varepsilon,\lambda}$-version of random Brouwer fixed point theorem is a special case of Theorem 2.1 (namely the $\mathcal{T}_{c}$-version of random Brouwer fixed point theorem). Hence, from now on, Theorem 2.1 is called random Brouwer fixed point theorem.
The focus of [26] lies in the proof of a noncompact Schauder fixed point theorem in random normed modules, where the authors merely presented a sketch proof of Theorem 2.1 based on the works [11, 28]. Considering the fundamental importance of Theorem 2.1, one of the purposes of this paper is to give a complete proof of Theorem 2.1 in order to make the work [26] self-contained. It is in the process of achieving the proof that many basic problems are treated thoroughly.
It is well known that the proof of the classical Brouwer fixed point theorem is not very easy and the simplest of all the classical proofs is by means of Sperner lemma [1, 33, 35, 48]. The classical Sperner lemma states that for any simplicial subdivision $\mathscr{S}$ of a simplex and any proper labeling function of $\mathscr{S}$ there are an odd number of completely labeled simplexes in $\mathscr{S}$. It is also well known that the most accessible proof of Sperner lemma was given by Kuhn in [35] via the graph-theoretic method. However, the proof of random Brouwer fixed point theorem involves more since $L^{0}(\mathcal{F},\mathbb{R}^{d})$ admits a more complicated topological and algebraic structure than $\mathbb{R}^{d}$, which requires the following three nontrivial tools. The first tool is Lemma 4.3 of [26], which states that a $\sigma$-stable mapping is $\mathcal{T}_{c}$-continuous iff it is random sequentially continuous. Since the topology $\mathcal{T}_{c}$ is too strong for us to conveniently treat the fixed point problem of the $\mathcal{T}_{c}$-continuous mappings, but random sequential continuity can be described in terms of $\mathcal{T}_{\varepsilon,\lambda}$, this lemma enables us to make full use of $\mathcal{T}_{\varepsilon,\lambda}$ to treat random Brouwer fixed point theorem. The second tool is the randomized version of the classical Bolzano-Weierstrass theorem [31], which states that an almost surely (briefly, a.s.) bounded sequence in $L^{0}(\mathcal{F},\mathbb{R}^{d})$ always admits an a.s. convergent randomized subsequence although it may admit no a.s. convergent subsequence. Since an a.s. bounded, $\mathcal{T}_{\varepsilon,\lambda}$-closed and $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$, is generally noncompact but random sequentially compact by the randomized Bolzano-Weierstrass theorem, then the randomized Bolzano-Weierstrass theorem can be used to overcome the difficulty due to noncompactness [11, 26, 28]. The third tool is a random version of the classical Sperner lemma. To prove a special case of the $\mathcal{T}_{\varepsilon,\lambda}$-version of random Brouwer fixed point theorem (see [11, Proposition 3.1]), the authors [11] first give a special case of random Sperner lemma (see the first part of [11, Theorem 2.3]), where the authors studied a special $L^{0}$-simplicial subdivision of an $L^{0}$-simplex (called the $L^{0}$-barycentric subdivision) in detail. The special random Sperner lemma provided in [11] was employed in all the works [11, 26, 28] on random Brouwer fixed point theorem. However, the proof of [11] on the special random Sperner lemma is not easily accessible to the reader, for example, the idea of the proof in [11] was very similar to Kuhn’s one [35], but the method in [11] was not by means of the graph-theoretic argumentation and thus rather complicated. Considering the fundamental importance of random Brouwer fixed point theorem, for example, it has been used in [26] for the proof of a noncompact Schauder fixed point theorem in random normed modules, and it will also be used to establish a noncompact $KKM$ principle in random locally convex modules so that a series of open problems posed in [26] will be solved in our forthcoming work, it is necessary to look for a very easily accessible proof of random Sperner lemma and to further give a complete proof of random Brouwer fixed point theorem. This leads us directly to systematically developing the theory of $L^{0}$-simplicial subdivisions of $L^{0}$-simplexes.
In fact, our approach to developing $L^{0}$-simplicial subdivisions is completely new, in particular our idea of defining an $L^{0}$-simplicial subdivision of an $L^{0}$-simplex is directly linking the $L^{0}$-simplicial subdivision to the corresponding counterpart simplicial subdivision of the corresponding counterpart simplex, see Definition 2.6. An advantage of Definition 2.6 enables us to establish a key representation theorem of a proper $L^{0}$-labeling function by usual proper labeling functions, see Theorem 3.5 of this paper for details. Consequently, by using the representation theorem we can smoothly establish a more general random Sperner lemma than the one given in [11], namely Theorem 2.9, in particular its proof is very easily accessible to the reader since this proof will be eventually finished by directly applying the classical Sperner lemma. Finally, combining Theorem 2.9 and the idea of the proof of Lemma 4.6 of [26], we can further give a complete proof of random Brouwer fixed point theorem.
Recently, we have known from [43, 44, 45] that Ponosov has been working on a fixed point theory for a class of local and continuous-in-probability operators with the aim of providing a tool for the study of stochastic differential equations. In particular, in [44] Ponosov proved a stochastic version of Brouwer fixed point theorem, which is stated as Proposition 2.4 in this paper and has played an essential role in the proofs of the two fixed point theorems — Theorems 5.1 and 5.2 of [45] that can be directly applied to the study of stochastic differential equations. In this paper we prove that Ponosov’s stochastic Brouwer fixed point theorem [44] is equivalent to the $\mathcal{T}_{\varepsilon,\lambda}$-version of random Brouwer fixed point theorem by combining Ponosov’s deep result [42] and the theory of complete random inner product modules. In particular, we further obtain a somewhat surprising conclusion again by means of Ponosov’s result [42]: a $\mathcal{T}_{\varepsilon,\lambda}$-continuous $\sigma$-stable mapping from a $\mathcal{T}_{\varepsilon,\lambda}$-closed $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ to $L^{0}(\mathcal{F},\mathbb{R}^{d})$ must be a.s. sequentially continuous (see Proposition 5.8 of this paper), which shows that Proposition 3.1 of [11] is equivalent to either of Proposition 2.3 and Proposition 2.4.
Since the classical Borsuk theorem [2] is of fundamental importance, which states that the unit sphere of an Euclidean space is not a retraction of the closed unit ball, and is also equivalent to classical Brouwer fixed point theorem. In [11], the $\mathcal{T}_{\varepsilon,\lambda}$-version of random Borsuk theorem was first proved, see Lemma 3.5 of [11], but the equivalence of the $\mathcal{T}_{\varepsilon,\lambda}$-version of random Borsuk theorem with the $\mathcal{T}_{\varepsilon,\lambda}$-version of random Brouwer fixed point theorem had been not yet given. In the end of this paper we not only prove the equivalence of the $\mathcal{T}_{\varepsilon,\lambda}$-version of random Borsuk theorem with the $\mathcal{T}_{\varepsilon,\lambda}$-version of random Brouwer fixed point theorem but also prove a general random Borsuk theorem and together with its equivalence with random Brouwer fixed point theorem.
This paper is organized as follows. In Section 2, we provide some prerequisites and a survey of different forms and special cases of random Brouwer fixed point theorem currently available. In Section 3, we establish a key representation theorem of a proper $L^{0}$-labeling function by usual proper labeling functions, namely Theorem 3.5. Based on Theorem 3.5, we prove Theorem 2.9, namely the random Sperner lemma. Next, as an application of Theorem 2.9, in Section 4, we prove Theorem 2.1. In Section 5, we prove the equivalence of Proposition 2.3 and Proposition 2.4. By the way, we also give some other interesting results in Section 5, which are closely related to the topic on random Brouwer fixed point theorem. Finally, in Section 6, we prove random Borsuk theorem — Theorem 6.4 is equivalent to random Brouwer fixed point theorem.
2 Random Brouwer fixed point theorem and Random Sperner lemm. Prerequisites
2.1 Different forms and special cases of random Brouwer fixed point theorem
To clearly present the main contents of this section, we first introduce some basic notation and terminologies.
Throughout this paper, $(\Omega,\mathcal{F},P)$ denotes a given probability space, $\mathbb{N}$ the set of positive integers and $\mathbb{K}$ the scalar field $\mathbb{R}$ of real numbers or $\mathbb{C}$ of complex numbers. We use $L^{0}(\mathcal{F},\mathbb{K})$ for the algebra of equivalence classes of $\mathbb{K}$-valued random variables on $(\Omega,\mathcal{F},P)$, and $L^{0}(\mathcal{F},\mathbb{R}^{d})$ for the linear space of equivalence classes of random vectors from $(\Omega,\mathcal{F},P)$ to $\mathbb{R}^{d}$. As usual, $L^{0}(\mathcal{F},\mathbb{R})$ is briefly denoted by $L^{0}(\mathcal{F})$. $L^{0}(\mathcal{F},\mathbb{R}^{d})$ forms a free $L^{0}(\mathcal{F})$-module of rank $d$. More generally, for any separable metric space $M$, $L^{0}(\mathcal{F},M)$ denotes the set of equivalence classes of $M$-valued $\mathcal{F}$-measurable functions on $(\Omega,\mathcal{F},P)$.
$\bar{L}^{0}(\mathcal{F})$ denotes the set of equivalence classes of extended real-valued random variables on $(\Omega,\mathcal{F},P)$. For any $\xi,\eta\in\bar{L}^{0}(\mathcal{F})$, we define $\xi\leq\eta$ iff $\xi^{0}(\omega)\leq\eta^{0}(\omega)$ for almost all $\omega$ in $\Omega$ (briefly, $\xi^{0}(\omega)\leq\eta^{0}(\omega)$ a.s.), where $\xi^{0}$ and $\eta^{0}$ are arbitrarily chosen representatives of $\xi$ and $\eta$, respectively, then $(\bar{L}^{0}(\mathcal{F}),\leq)$ forms a complete lattice (see [12] for details).
As usual, $\xi<\eta$ means $\xi\leq\eta$ and $\xi\neq\eta$, whereas, for any $A\in\mathcal{F}$, $\xi<\eta$ on $A$ means $\xi^{0}(\omega)<\eta^{0}(\omega)$ for almost all $\omega\in A$, where $\xi^{0}$ and $\eta^{0}$ are again arbitrary representatives of $\xi$ and $\eta$, respectively. Furthermore, we denote $L_{+}^{0}(\mathcal{F}):=\{\xi\in L^{0}(\mathcal{F}):\xi\geq 0\}$ and $L_{++}^{0}(\mathcal{F}):=\{\xi\in L^{0}(\mathcal{F}):\xi>0~{}on~{}\Omega\}$. For any $A\in\mathcal{F}$, we use $\tilde{I}_{A}$ for the equivalence class of $I_{A}$, where $I_{A}$ is the characteristic function of $A$ (namely $I_{A}(\omega)=1$ if $\omega\in A$ and $I_{A}(\omega)=0$ otherwise).
For any $\xi,\eta\in\bar{L}^{0}(\mathcal{F})$, we use $(\xi=\eta)$ for the equivalence class of $(\xi^{0}=\eta^{0})$, where $(\xi^{0}=\eta^{0}):=\{\omega\in\Omega:\xi^{0}(\omega)=\eta^{0}(\omega)\}$, and $\xi^{0}$ and $\eta^{0}$ are arbitrarily chosen representatives of $\xi$ and $\eta$, respectively. Similarly, one can understand $(\xi<\eta)$ and $(\xi\leq\eta)$.
[…]
A subset $G$ of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is said to be
  • a.s. bounded if $\bigvee\{\|x\|:x\in G\}\in L_{+}^{0}(\mathcal{F})$.
  • $L^{0}$-convex if $\lambda x+(1-\lambda)y\in G$ for any $x,y\in G$ and any $\lambda\in L_{+}^{0}(\mathcal{F})$ with $\lambda\leq 1$.
  • $\sigma$-stable if $\sum_{n=1}^{\infty}\tilde{I}_{A_{n}}x_{n}\in G$ for any sequence $\{x_{n},n\in\mathbb{N}\}$ in $G$ and any countable partition $\{A_{n},n\in\mathbb{N}\}$ of $\Omega$ to $\mathcal{F}$ (namely each $A_{n}\in\mathcal{F},A_{i}\cap A_{j}=\emptyset$ for any $i\neq j$, and $\bigcup_{n=1}^{\infty}A_{n}=\Omega$). Here, $\sum_{n=1}^{\infty}\tilde{I}_{A_{n}}x_{n}$ is well defined, namely it has a natural explanation: arbitrarily choosing a representative $x_{n}^{0}$ of $x_{n}$ for each $n\in\mathbb{N}$, the function $\sum_{n=1}^{\infty}I_{A_{n}}x_{n}^{0}$ defined by $(\sum_{n=1}^{\infty}I_{A_{n}}x_{n}^{0})(\omega)=x_{n}^{0}(\omega)$ when $\omega\in A_{n}$ for some $n\in\mathbb{N}$, is exactly a representative of $\sum_{n=1}^{\infty}\tilde{I}_{A_{n}}x_{n}$.
A mapping $f:G\rightarrow G$ is said to be
  • $\sigma$-stable if $G$ is $\sigma$-stable and $f(\sum_{n=1}^{\infty}\tilde{I}_{A_{n}}x_{n})=\sum_{n=1}^{\infty}\tilde{I}_{A_{n}}f(x_{n})$ for any sequence $\{x_{n},n\in\mathbb{N}\}$ in $G$ and any countable partition $\{A_{n},n\in\mathbb{N}\}$ of $\Omega$ to $\mathcal{F}$.
  • $\mathcal{T}_{\varepsilon,\lambda}$-continuous if $f$ is a continuous mapping from $(G,\mathcal{T}_{\varepsilon,\lambda})$ to $(G,\mathcal{T}_{\varepsilon,\lambda})$.
  • $\mathcal{T}_{c}$-continuous if $f$ is a continuous mapping from $(G,\mathcal{T}_{c})$ to $(G,\mathcal{T}_{c})$.
  • a.s. sequentially continuous if for any $x\in G$ and any sequence $\{x_{n},n\in\mathbb{N}\}$ in $G$ such that $\{\|x_{n}-x\|,n\in\mathbb{N}\}$ converges a.s. to 00, it holds that $\{\|f(x_{n})-f(x)\|,n\in\mathbb{N}\}$ converges a.s. to 00.
  • random sequentially continuous if $G$ is $\sigma$-stable and for any $x\in G$ and any sequence $\{x_{n},n\in\mathbb{N}\}$ in $G$ convergent in $\mathcal{T}_{\varepsilon,\lambda}$ to $x$ there exists a random subsequence $\{x_{n_{k}},k\in\mathbb{N}\}$ of $\{x_{n},n\in\mathbb{N}\}$ such that $\{f(x_{n_{k}}),k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $f(x)$. Here, $\{x_{n_{k}},k\in\mathbb{N}\}$ is a random subsequence of $\{x_{n},n\in\mathbb{N}\}$, we means $\{n_{k},k\in\mathbb{N}\}$ is a sequence of equivalence classes of positive-integer-valued random variables such that the following two conditions are satisfied:
    • (1)
      $n_{k}<n_{k+1}$ on $\Omega$ for any $k\in\mathbb{N}$;
    • (2)
      $x_{n_{k}}=\sum_{l=1}^{\infty}I_{(n_{k}=l)}x_{l}$ for any $k\in\mathbb{N}$, where $n_{k}^{0}$ is an arbitrarily chosen representative of $n_{k}$, $(n_{k}=l)$ is the equivalence class of $A_{k,l}:=\{\omega\in\Omega:n_{k}^{0}(\omega)=l\}$ and $I_{(n_{k}=l)}=\tilde{I}_{A_{k,l}}$.
One of the main purposes of this paper is to provide a complete proof of the following random Brouwer fixed point theorem — Theorem 2.1, which was given a sketch proof in [26] (see Lemma 4.6 of [26]).
Theorem 2.1.
Let $G$ be a $\sigma$-stable, a.s. bounded, $\mathcal{T}_{c}$-closed and $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ and $f:G\rightarrow G$ be a $\sigma$-stable $\mathcal{T}_{c}$-continuous mapping. Then $f$ has a fixed point.
Since $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is a $\mathcal{T}_{\varepsilon,\lambda}$-complete random normed module, every $\mathcal{T}_{\varepsilon,\lambda}$-closed and $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ must be $\sigma$-stable, and Lemma 4.3 of [26] shows that a $\sigma$-stable mapping $f$ is $\mathcal{T}_{c}$-continuous iff $f$ is random sequentially continuous. It is also known from [23] that a $\sigma$-stable subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is $\mathcal{T}_{\varepsilon,\lambda}$-closed iff it is $\mathcal{T}_{c}$-closed, so Proposition 2.2 below, which was first given as Lemma 4.6 of [26], is an equivalent formulation of Theorem 2.1.
Proposition 2.2.
Let $G$ be an a.s. bounded, $\mathcal{T}_{\varepsilon,\lambda}$-closed and $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ and $f:G\rightarrow G$ be a $\sigma$-stable random sequentially continuous mapping. Then $f$ has a fixed point.
It is obvious that a $\mathcal{T}_{\varepsilon,\lambda}$-continuous mapping defined on a $\sigma$-stable set is random sequentially continuous, Proposition 2.3 below, which first occurred in [28, Proposition 4.2] is a special case of Proposition 2.2.
Proposition 2.3.
Let $G$ be an a.s. bounded, $\mathcal{T}_{\varepsilon,\lambda}$-closed and $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ and $f:G\rightarrow G$ be a $\sigma$-stable $\mathcal{T}_{\varepsilon,\lambda}$-continuous mapping. Then $f$ has a fixed point.
Let $X$ and $Y$ be two separable metric spaces, it is well known that $L^{0}(\mathcal{F},X)$ and $L^{0}(\mathcal{F},Y)$ are metrizable spaces when they are endowed with the topologies of convergence in probability, respectively. Further, let $G$ be a nonempty subset of $L^{0}(\mathcal{F},X)$, a mapping $h:G\rightarrow L^{0}(\mathcal{F},Y)$ is said to be local if $x^{0}(\omega)=y^{0}(\omega)$ a.s. on $A$ implies $h(x)^{0}(\omega)=h(y)^{0}(\omega)$ a.s. on $A$ for any $x$ and $y$ in $G$ and any $A\in\mathcal{F}$, where $x^{0}$, $y^{0}$, $h(x)^{0}$ and $h(y)^{0}$ are any given representatives of $x$, $y$, $h(x)$ and $h(y)$, respectively. Finally, $h$ is said to be continuous in probability if $h$ is a continuous mapping from $G$ to $L^{0}(\mathcal{F},Y)$ when $G$ and $L^{0}(\mathcal{F},Y)$ are endowed with the topologies of convergence in probability. In fact, a mapping is continuous in probability iff it is $\mathcal{T}_{\varepsilon,\lambda}$-continuous.
As usual, $2^{\mathbb{R}^{d}}$ denotes the family of subsets of $\mathbb{R}^{d}$. Let $U:\Omega\rightarrow 2^{\mathbb{R}^{d}}$ be a set-valued mapping satisfying the following two conditions: (1) $U(\omega)$ is a nonempty bounded closed convex subset of $\mathbb{R}^{d}$ for each $\omega\in\Omega$; (2) $Gr(U):=\{(\omega,x)\in\Omega\times\mathbb{R}^{d}:x\in U(\omega)\}\in\mathcal{F}\otimes Bor(\mathbb{R}^{d})$, namely $U$ has a measurable graph, where $Bor(\mathbb{R}^{d})$ stands for the Borel $\sigma$-algebra of $\mathbb{R}^{d}$.
Further, for $U$ as above, we always denote by $G_{U}$ the set of equivalence classes of almost surely measurable selections of $U$, namely $G_{U}:=\{x\in L^{0}(\mathcal{F},\mathbb{R}^{d}):x$ has a representative $x^{0}$ such that $x^{0}(\omega)\in U(\omega)~{}a.s.\}$.
Then it is not very difficult to verify that $G_{U}$ is an a.s. bounded $\mathcal{T}_{\varepsilon,\lambda}$-closed $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ (see Lemma 5.3 of this paper), so $G_{U}$ is also $\sigma$-stable and it is easy to check that a mapping $h:G_{U}\rightarrow G_{U}$ is local iff $h$ is $\sigma$-stable (see Lemma 5.1 of this paper).
Proposition 2.4 below, which was called the stochastic Brouwer fixed point theorem in [44] and has played an essential role in the proofs of Theorems 5.1 and 5.2 of [45]. It is obvious that Proposition 2.4 is a special case of Proposition 2.3. The original proof of Proposition 2.4 was based on Ponosov’s solution of the Nemytskii conjecture [42] and a measurable selection theorem from [51]. In this paper, we can even prove that Proposition 2.4 is equivalent to Proposition 2.3 by means of Ponosov’s result [42]. In particular, we can also obtain a somewhat surprising conclusion again by means of Ponosov’s result [42]: a $\mathcal{T}_{\varepsilon,\lambda}$-continuous $\sigma$-stable mapping from a $\mathcal{T}_{\varepsilon,\lambda}$-closed $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ to $L^{0}(\mathcal{F},\mathbb{R}^{d})$ must be a.s. sequentially continuous, see Proposition 5.8 of this paper.
Proposition 2.4 is a slightly more general formulation of Theorem 2.1 of [44] (see also Lemma B.4 of [45]) in that we remove the assumption on completeness of $(\Omega,\mathcal{F},P)$ (otherwise, we first consider $(\Omega,\hat{\mathcal{F}},P)$, which is the completion of $(\Omega,\mathcal{F},P)$, and obtain a fixed point $x$ of $h$ as in [44] and $x$ has a representative $x^{0}$ such that $x^{0}(\omega)\in U(\omega)$ a.s. and $x^{0}$ is $\hat{\mathcal{F}}$-measurable, whereas $x^{0}$ is equal a.s. to an $\mathcal{F}$-measurable function $\hat{x}^{0}$, then the equivalence class $\hat{x}$ of $\hat{x}^{0}$ is exactly our desired fixed point).
Proposition 2.4.
Let $U:\Omega\rightarrow 2^{\mathbb{R}^{d}}$ be a set-valued function with a measurable graph such that $U(\omega)$ is a nonempty bounded closed convex subset of $\mathbb{R}^{d}$ for each $\omega\in\Omega$. Then every local and continuous-in-probability operator from $G_{U}$ to itself has a fixed point.
A very special case of Proposition 2.3 first appeared as Proposition 3.1 in [11], which states that every $\sigma$-stable a.s. sequentially continuous mapping from an a.s. bounded a.s. sequentially closed $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ to itself has a fixed point, where a subset is said to be a.s. sequentially closed if it contains all a.s. limits of sequences in it. Proposition 2.3 is a generalization of Proposition 3.1 of [11] since a.s. sequentially closed set is exactly $\mathcal{T}_{\varepsilon,\lambda}$-closed and a.s. sequential continuity obviously implies $\mathcal{T}_{\varepsilon,\lambda}$-continuity. It is interesting to notice that Proposition 5.8 implies the equivalence of Proposition 3.1 of [11] with either of Proposition 2.3 and Proposition 2.4.
We would also like to point out that Proposition 2.2 (namely Theorem 2.1) and Proposition 2.3 are both best possible, which is explained as follows. It is well known that $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is a metrizable linear topological space under $\mathcal{T}_{\varepsilon,\lambda}$ (namely the topology of convergence in probability), there is, naturally, the notion of boundedness in the sense of the linear topology $\mathcal{T}_{\varepsilon,\lambda}$, denoted by the $\mathcal{T}_{\varepsilon,\lambda}$-boundedness, it is easy to see that the a.s. boundedness implies the $\mathcal{T}_{\varepsilon,\lambda}$-boundedness, and we have known from Proposition 16 of [24] that these two kinds of boundedness coincide for a $\sigma$-stable subset. Likewise, the $L^{0}$-convexity implies the usual convexity, and we have also known from Lemma 2.8 of [49] that these two kinds of convexity coincide for a $\mathcal{T}_{\varepsilon,\lambda}$-closed and $\sigma$-stable subset. On the one hand, an a.s. bounded $\mathcal{T}_{\varepsilon,\lambda}$-closed $L^{0}$-convex subset $G$ of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is generally noncompact, if the hypothesis on the $\sigma$-stability of $f$ in Proposition 2.3 is removed, then Proposition 2.3 does not hold, which comes from a classical result due to Dobrowolski and Marciszewski [10] (see also [26, Section 2.1]). On the other hand, an example provided by Ponosov in [44] shows that Proposition 2.4 (namely Proposition 2.3) does not hold if $G_{U}$ is replaced with a generic $\mathcal{T}_{\varepsilon,\lambda}$-closed, $\mathcal{T}_{\varepsilon,\lambda}$-bounded and convex set $G$ (namely $G$ is not $L^{0}$-convex). Therefore, they are all best possible.
2.2 A general $L^{0}$-simplicial subdivision and random Sperner lemma
In this paper, we present a general theory of $L^{0}$-simplicial subdivisions of $L^{0}$-simplexes and give a thorough treatment of random Sperner lemma by grasping the inherent connection between $L^{0}$-simplexes and classical simplexes, so we can smoothly present a good definition of an $L^{0}$-simplicial subdivision of an $L^{0}$-simplex (see Definition 2.6 below) and establish a general random Sperner lemma (see Theorem 2.9 below). To survey these main results, let us first introduce some necessary notation and prerequisites.
In the sequel of this paper, a left module over the algebra $L^{0}(\mathcal{F},\mathbb{K})$ is simply said to be an $L^{0}$-module.
Since $L^{0}$-labeling functions take values in $L^{0}(\mathcal{F},\{1,2,\cdots,n\})$ (see Definition 2.7 below), to remove any possible ambiguities by strictly distinguishing a measurable set and its equivalence class, we employ the language from measure algebras. For this, we introduce the convention from measure algebras as follows.
For any $A,B\in\mathcal{F}$, we will always use the corresponding lowercase letters $a$ and $b$ for the equivalence classes $[A]$ and $[B]$ (two elements $C$ and $D$ in $\mathcal{F}$ are said to be equivalent if $P[(C\setminus D)\cup(D\setminus C)]=0$), respectively. Define $1=[\Omega]$, $0=[\emptyset]$, $a\wedge b=[A\cap B]$, $a\vee b=[A\cup B]$ and $a^{c}=[A^{c}]$, where $A^{c}$ denotes the complement of $A$. Let $B_{\mathcal{F}}=\{a=[A]:A\in\mathcal{F}\}$, then $(B_{\mathcal{F}},\wedge,\vee,^{c},0,1)$ is a complete Boolean algebra, namely a complete complemented distributive lattice (see [34] for details). Specially, $B_{\mathcal{F}}$ is called the measure algebra associated with $(\Omega,\mathcal{F},P)$.
As usual, for any $a,b\in B_{\mathcal{F}}$, $a>b$ means $a\geq b$ and $a\neq b$. A subset $\{a_{i}:i\in I\}$ of $B_{\mathcal{F}}$ is called a partition of unity if $\bigvee_{i\in I}a_{i}=1$ and $a_{i}\wedge a_{j}=0$ for any $i,j\in I$ with $i\neq j$. The collection of all such partitions is denoted by $p(1)$. For any $\{a_{j}:j\in J\}\in p(1)$, it is clear that the set $\{j\in J:a_{j}>0\}$ is at most countable. Moreover, $\{a_{n},n\in\mathbb{N}\}\in p(1)$ iff there exists a countable partition $\{A_{n},n\in\mathbb{N}\}$ of $\Omega$ to $\mathcal{F}$ such that $a_{n}=[A_{n}]$ for each $n\in\mathbb{N}$.
From now on, for any $a=[A]\in B_{\mathcal{F}}$, we also use $I_{a}$ to denote $\tilde{I}_{A}$. For a subset $H$ of a complete lattice (e.g., $B_{\mathcal{F}}$), we use $\bigvee H$ and $\bigwedge H$ to denote the supremum and infimum of $H$, respectively.
Recall that an $L^{0}$-module $E$ is said to be regular if $E$ has the following property: for any given two elements $x$ and $y$ in $E$, if there exists $\{a_{n},n\in\mathbb{N}\}\in p(1)$ such that $I_{a_{n}}x=I_{a_{n}}y$ for each $n\in\mathbb{N}$, then $x=y$.
In the remainder of this paper, all the $L^{0}$-modules under consideration are assumed to be regular. This restriction is not excessive, since all random normed modules (for example, $L^{0}(\mathcal{F},\mathbb{R}^{d})$) are regular.
Since our future target is to develop random functional analysis and its applications to dynamic Nash equilibrium theory and nonsmooth differential geometry on metric measure spaces (see [4, 5, 6, 16, 17, 36, 37] for recent advances), random normed modules and random locally convex modules will play an active role and they are all more general $L^{0}$-modules than $L^{0}(\mathcal{F},\mathbb{R}^{d})$, and thus we state the results of this section in the $L^{0}$-modules for possible convenience in the future applications.
In the special $L^{0}$-module $L^{0}(\mathcal{F},\mathbb{R}^{d})$, $\sum_{n=1}^{\infty}I_{a_{n}}x_{n}$ is always well defined for any sequence $\{x_{n},n\in\mathbb{N}\}$ and any $\{a_{n},n\in\mathbb{N}\}\in p(1)$. However, for a generic $L^{0}$-module $E$, $\sum_{n=1}^{\infty}I_{a_{n}}x_{n}$ may be meaningless for any sequence $\{x_{n},n\in\mathbb{N}\}$ for any sequence $\{x_{n},n\in\mathbb{N}\}$ in $E$ and any $\{a_{n},n\in\mathbb{N}\}\in p(1)$, but once a nonempty subset $G$ of $E$ is $\sigma$-stable, $\sum_{n=1}^{\infty}I_{a_{n}}x_{n}$ is well defined for any sequence $\{x_{n},n\in\mathbb{N}\}$ in $G$ and any $\{a_{n},n\in\mathbb{N}\}\in p(1)$. The notion of a $\sigma$-stable set was introduced in [23] (see also (4) of Definition 2.5 below) in order to make sense the countable concatenation such as $\sum_{n=1}^{\infty}I_{a_{n}}x_{n}$.
Definition 2.5 ([23, 26]).
Let $E$ be an $L^{0}$-module and $G$ be a nonempty subset of $E$.
  1. (1)
    $G$ is said to be $L^{0}$-convex if $\lambda x+(1-\lambda)y\in G$ for any $x,y\in G$ and any $\lambda\in L_{+}^{0}(\mathcal{F})$ with $\lambda\leq 1$. Generally, for any subset $G$, $Conv_{L^{0}}(G):=\{\sum^{n}_{i=1}\xi_{i}x_{i}:n\in\mathbb{N},\{x_{1},x_{2},\cdots,x_{n}\}\subseteq G~{}\text{and}~{}\{\xi_{1},\xi_{2},\cdots,\xi_{n}\}\subseteq L^{0}_{+}(\mathcal{F})~{}\text{such that}~{}\sum^{n}_{i=1}\xi_{i}=1\}$ is called the $L^{0}$-convex hull of $G$. Clearly, $G$ is $L^{0}$-convex if $Conv_{L^{0}}(G)=G$.
  2. (2)
    A finite subset $\{x_{1},x_{2},\cdots,x_{n}\}$ of $G$ is said to be $L^{0}$-linearly independent if, for any $\{\xi_{1},\xi_{2},\cdots,\xi_{n}\}\subseteq L^{0}(\mathcal{F},\mathbb{K})$ satisfying $\sum^{n}_{i=1}\xi_{i}x_{i}=0$, it always holds that $\xi_{i}=0$ for each $i=1\sim n$. Further, $G$ is said to be $L^{0}$-linearly independent if any finite nonempty subset of $G$ is $L^{0}$-linearly independent, in which case the $L^{0}$-submodule generated by $G$ is said to be $L^{0}$-free.
  3. (3)
    A finite subset $\{x_{1},x_{2},\cdots,x_{n}\}$ of $E$ is said to be $L^{0}$-affinely independent if either $n=1$ or if $n>1$ and $\{x_{i}-x_{1}:i=2\sim n\}$ is $L^{0}$-linearly independent, in which case $Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ is called an $(n-1)$-dimensional $L^{0}$-simplex with vertexes $x_{1},x_{2},\cdots,x_{n}$ and $Conv_{L^{0}}(\{x_{i}:i\in I\})$ is called an $(|I|-1)$-dimensional $L^{0}$-face of $Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ for any nonempty subset $I$ of $\{1,2,\cdots,n\}$, where $|I|$ denotes the cardinality of $I$.
  4. (4)
    $G$ is said to be finitely stable if $I_{a}x+I_{a^{c}}y\in G$ for any $x,y\in G$ and any $a\in B_{\mathcal{F}}$. $G$ is said to be $\sigma$-stable (or to have the countable concatenation property in terms of [23]) if, for each sequence $\{x_{n},n\in\mathbb{N}\}$ in $G$ and each $\{a_{n},n\in\mathbb{N}\}\in p(1)$, there exists some $x\in G$ such that $I_{a_{n}}x=I_{a_{n}}x_{n}$ for each $n\in\mathbb{N}$, such an $x$ is unique since $E$ is assumed to be regular, usually denoted by $\sum_{n=1}^{\infty}I_{a_{n}}x_{n}$, called the countable concatenation of $\{x_{n},n\in\mathbb{N}\}$ along $\{a_{n},n\in\mathbb{N}\}$. By the way, if $G$ is $\sigma$-stable and $H$ is a nonempty subset of $G$, then $\sigma(H):=\{\sum_{n=1}^{\infty}I_{a_{n}}h_{n}:\{h_{n},n\in\mathbb{N}\}$ is a sequence in $H$ and $\{a_{n},n\in\mathbb{N}\}\in p(1)\}$ is called the $\sigma$-stable hull of $H$.
For any $(n-1)$-dimensional $L^{0}$-simplex $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ in an $L^{0}$-module $E$, as shown by Guo et al. in [26], $S$ is $\sigma$-stable and each element $x\in S$ can be uniquely represented as $x=\sum^{n}_{i=1}\xi_{i}x_{i}$, where $\xi_{i}\in L^{0}_{+}(\mathcal{F})$ for any $i\in\{1,2,\cdots,n\}$ and are such that $\sum^{n}_{i=1}\xi_{i}=1$. Moreover, since $E$ is also a linear space over $\mathbb{K}$ and the $L^{0}$-affine independence of $\{x_{1},x_{2},\cdots,x_{n}\}$ implies their affine independence, then $S^{\prime}=Conv(\{x_{1},x_{2},\cdots,x_{n}\})$ is a classical $(n-1)$-dimensional simplex in the linear space $E$, where $Conv(G)$ denotes the usual convex hull of a nonempty subset $G$ in a linear space. In this case, we call $S^{\prime}$ the counterpart simplex of $S$. Throughout this paper, whenever we refer to an $L^{0}$-simplex $S$, we will always denote its counterpart simplex by $S^{\prime}$.
Recall from [1] that for any $(n-1)$-dimensional simplex $,S^{\prime}=Conv(\{x_{1},x_{2},$$\cdots,x_{n}\})$ in a linear space, a finite family $\mathscr{S}^{\prime}$ of $(n-1)$-dimensional simplexes is called a simplicial subdivision of $S^{\prime}$ if
  • (1)
    $S^{\prime}=\bigcup\mathscr{S}^{\prime}$;
  • (2)
    for any $S^{\prime}_{1},S^{\prime}_{2}\in\mathscr{S}^{\prime}$, $S^{\prime}_{1}\cap S^{\prime}_{2}$ is either empty or a common face of $S^{\prime}_{1}$ and $S^{\prime}_{2}$.
As a random generalization, we introduce the notion of an $L^{0}$-simplicial subdivision of an $L^{0}$-simplex in Definition 2.6 below.
Definition 2.6.
Let $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ be an $(n-1)$-dimensional $L^{0}$-simplex in an $L^{0}$-module $E$. A finite family $\mathscr{S}$ of $(n-1)$-dimensional $L^{0}$-simplexes is called an $L^{0}$-simplicial subdivision of $S$ if
  1. (1)
    $S=\sigma\{\bigcup\mathscr{S}\}$;
  2. (2)
    for any $S_{1},S_{2}\in\mathscr{S}$, $S_{1}\cap S_{2}$ is either empty or a common $L^{0}$-face of $S_{1}$ and $S_{2}$;
  3. (3)
    $\mathscr{S}^{\prime}=\{G^{\prime}:G\in\mathscr{S}\}$ is a simplicial subdivision of $S^{\prime}$.
Let $\mathscr{S}$ and $\mathscr{S}^{\prime}$ be the same as in Definition 2.6, then we call $\mathscr{S}^{\prime}$ the counterpart simplicial subdivision of $\mathscr{S}$. Throughout this paper, whenever we refer to an $L^{0}$-simplicial subdivision $\mathscr{S}$, we will always denote its counterpart simplicial subdivision by $\mathscr{S}^{\prime}$. $L^{0}$-simplexes are much more complicated objects than usual simplexes, we encounter lots of challenges in the initial study of them, these force us to introduce Definition 2.6 in connection with the classical simplicial subdivision so that we can make full use of the theory of classical simplicial subdivisions to give a thorough treatment of $L^{0}$-simplicial subdivisions of $L^{0}$-simplexes.
The notion of $L^{0}$-extreme points was introduced and studied by Guo et al. in [26]. For two nonempty subsets $G$ and $H$ of an $L^{0}$-module (resp., a linear sapce) such that $H\subseteq G$, $H$ is called an $L^{0}$-extreme subset (resp., an extreme subset) of $G$ if both $x$ and $y$ belong to $H$ whenever $x$ and $y$ belong to $G$ and there exists $\lambda\in L^{0}(\mathcal{F},(0,1))$ (resp., $\lambda\in(0,1)$) such that $\lambda x+(1-\lambda)y\in H$. Further, if $H=\{h\}$ is a singleton and an $L^{0}$-extreme subset (resp., an extreme subset) of $G$, then $h$ is called an $L^{0}$-extreme point (resp., an extreme point) of $G$. We will always denote by $ext_{L^{0}}(G)$ (resp., $ext(G)$) the set of $L^{0}$-extreme points (resp., extreme points) of $G$. For any family $\mathscr{G}$ of nonempty subsets of an $L^{0}$-module (resp., a linear space), the set of $L^{0}$-extreme points (resp., extreme points) of $\mathscr{G}$ is defined by $ext_{L^{0}}(\mathscr{G}):=\sigma[\bigcup_{G\in\mathscr{G}}ext_{L^{0}}(G)]$ (resp., $ext(\mathscr{G}):=\bigcup_{G\in\mathscr{G}}ext(G)$).
For any $(n-1)$-dimensional simplex $S^{\prime}=Conv(\{x_{1},x_{2},\cdots,x_{n}\})$ in a linear space and any simplicial subdivision $\mathscr{S}^{\prime}$ of $S^{\prime}$, a mapping $\psi:ext(\mathscr{S}^{\prime})\rightarrow\{1,2,\cdots,n\}$ is called a proper labeling function of $\mathscr{S}^{\prime}$ (see [1]) if
$\psi(y)\in\chi(y),\forall y\in ext(\mathscr{S}^{\prime}),$
where $\chi(y):=\{i:\alpha_{i}>0\}$ for any $y=\sum_{i=1}^{n}\alpha_{i}x_{i}$. In fact, the set of proper labeling functions is exactly $\Pi_{y\in ext(\mathscr{S}^{\prime})}\chi(y)$.
A simplex $G^{\prime}=Conv(\{y_{1},y_{2},\cdots,y_{n}\})$ with $y_{j}\in ext(\mathscr{S}^{\prime})$ for each $j\in\{1,2,\cdots,n\}$ is said to be completely labeled by $\psi$ if $\psi$ is a proper labeling function of $\mathscr{S}^{\prime}$ and
$\{\psi(y_{j}):j=1\sim n\}=\{1,2,\cdots,n\}.$
Let $E_{1}$ and $E_{2}$ be two $L^{0}$-modules, and let $G_{1}\subseteq E_{1}$ and $G_{2}\subseteq E_{2}$ be two $\sigma$-stable sets. A mapping $f:G_{1}\rightarrow G_{2}$ is said to be $\sigma$-stable if
$f(\sum^{\infty}_{n=1}I_{a_{n}}x_{n})=\sum^{\infty}_{n=1}I_{a_{n}}f(x_{n})$
for any $\{a_{n},n\in\mathbb{N}\}\in p(1)$ and any sequence $\{x_{n},n\in\mathbb{N}\}$ in $G_{1}$.
In [11, Definition 2.1], Drapeau et al. introduced the notion of a proper $L^{0}$-labeling function for the $L^{0}$-barycentric subdivision of an $L^{0}$-simplex in $L^{0}(\mathcal{F},\mathbb{R}^{d})$. Definition 2.7 below is a generalization of [11, Definition 2.1]. But we adopt a strengthened notion of a completely labeled $L^{0}$-simplex since such a notion is both easy to understand and such a completely labeled $L^{0}$-simplex indeed exists (see Theorem 2.9). In fact, according to Definition 2.7, $\phi(y_{j})$ is an equivalence class of a constant valued random variable for each $j\in\{1,2,\cdots,n\}$ for a completely labeled $L^{0}$-simplex $G=Conv_{L^{0}}(\{y_{1},y_{2},\cdots,y_{n}\})$, and hence our notion of a completely labeled $L^{0}$-simplex is completely similar to the notion of a completely labeled simplex.
Definition 2.7.
Let $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ be an $(n-1)$-dimensional $L^{0}$-simplex in an $L^{0}$-module $E$, $\mathscr{S}$ an $L^{0}$-simplicial subdivision of $S$ and $\phi:ext_{L^{0}}(\mathscr{S})\rightarrow L^{0}(\mathcal{F},\{1,2,\cdots,n\})$ a mapping. $\phi$ is called an $L^{0}$-labeling function of $\mathscr{S}$ if $\phi$ is $\sigma$-stable. An $L^{0}$-labeling function $\phi$ of $\mathscr{S}$ is said to be proper if
$(\phi(y)=i)\wedge(\lambda_{i}=0)=0~{}\text{for~{}any}~{}y\in ext_{L^{0}}(\mathscr{S})~{}\text{and~{}any}~{}i\in\{1,2,\cdots,n\},$
where $y=\sum_{i=1}^{n}\lambda_{i}x_{i}$. An $L^{0}$-simplex $G=Conv_{L^{0}}(\{y_{1},y_{2},\cdots,y_{n}\})$ with $y_{j}\in ext_{L^{0}}(\mathscr{S})$ for each $j\in\{1,2,\cdots,n\}$ is said to be completely labeled by $\phi$ if $\phi$ is a proper $L^{0}$-labeling function of $\mathscr{S}$ and
$\{\phi(y_{j}):j=1\sim n\}=\{1,2,\cdots,n\}.$
The classical Sperner lemma, originally established by Sperner in [48] was formulated for simplexes in $\mathbb{R}^{n}$. Let $S^{\prime}=Conv(\{x_{1},x_{2},\cdots,x_{n}\})$ be an $(n-1)$-dimensional simplex in a linear space $L$ and $L^{n}=Span(\{x_{1},x_{2},\cdots,x_{n}\})$ be the linear subspace of $L$ generated by $\{x_{1},x_{2},\cdots,x_{n}\}$, then $L^{n}$ is at most $n$-dimensional and the classical Sperner lemma is only related to $L^{n}$ and is independent of $L$, so one can readily see the following version of Sperner lemma for simplexes in a general linear space.
Lemma 2.8.
Let $S^{\prime}=Conv(\{x_{1},x_{2},\cdots,x_{n}\})$ be an $(n-1)$-dimensional simplex in a linear space, $\mathscr{S}^{\prime}$ a simplicial subdivision of $S^{\prime}$ and $\psi$ a proper labeling function of $\mathscr{S}^{\prime}$. Then there are an odd number of completely labeled simplexes in $\mathscr{S}^{\prime}$
For any given sequence $\{G_{n},n\in\mathbb{N}\}$ of nonempty subsets of a $\sigma$-stable set $G$ and any given $\{a_{n},n\in\mathbb{N}\}\in p(1)$,
$\sum_{n=1}^{\infty}I_{a_{n}}G_{n}:=\{\sum_{n=1}^{\infty}I_{a_{n}}x_{n}:x_{n}\in G_{n}~{}\text{for~{}each}~{}n\in\mathbb{N}\}$
is called the countable concatenation of $\{G_{n},n\in\mathbb{N}\}$ along $\{a_{n},n\in\mathbb{N}\}$.
Theorem 2.9 below can be aptly called the random Sperner lemma.
Theorem 2.9.
Let $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ be an $(n-1)$-dimensional $L^{0}$-simplex in an $L^{0}$-module $E$, $\mathscr{S}$ an $L^{0}$-simplicial subdivision of $S$ and $\phi$ a proper $L^{0}$-labeling function of $\mathscr{S}$. Then there exist a finite family $\{S_{m}:=Conv_{L^{0}}(\{y_{1}^{m},y_{2}^{m},\cdots,y_{n}^{m}\}),m=1\sim M\}$ of $(n-1)$-dimensional $L^{0}$-simplexes in $\mathscr{S}$ and $\{a_{m},m=1\sim M\}\in p(1)$ such that
$\sum_{m=1}^{M}I_{a_{m}}S_{m}=Conv_{L^{0}}(\{\sum_{m=1}^{M}I_{a_{m}}y_{1}^{m},\sum_{m=1}^{M}I_{a_{m}}y_{2}^{m},\cdots,\sum_{m=1}^{M}I_{a_{m}}y_{n}^{m}\})$
is an $(n-1)$-dimensional $L^{0}$-simplex completely labeled by $\phi$ and satisfies that
$\phi(\sum_{m=1}^{M}I_{a_{m}}y_{j}^{m})=j~{}\text{for~{}any}~{}j\in\{1,2,\cdots,n\}.$
In Theorem 2.9, $\sum_{m=1}^{M}I_{a_{m}}S_{m}$ is an $(n-1)$-dimensional $L^{0}$-simplex contained in $S$ such that each vertex of $\sum_{m=1}^{M}I_{a_{m}}S_{m}$ belongs to $ext_{L^{0}}(\mathscr{S})$ since both $S$ and $ext_{L^{0}}(\mathscr{S})$ are $\sigma$-stable. Besides, $\sum_{m=1}^{M}I_{a_{m}}S_{m}$ is not necessarily in $\mathscr{S}$ but merely a finite concatenation of the finite subfamily $\{S_{m}:m=1\sim M\}$ of $\mathscr{S}$, which is a significant difference between Theorem 2.9 and the classical Sperner lemma (namely Lemma 2.8).
3 Proof of Theorem 2.9
We first give the following two Examples — Examples 3.1 and 3.2. The first one shows the $L^{0}$-barycentric subdivision introduced in [11, Definition 1.10] is a special case of our Definition 2.6. The second one shows our Definition 2.6 does not just include the $L^{0}$-barycentric subdivision.
Example 3.1.
Let $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ be an $(n-1)$-dimensional $L^{0}$-simplex in an $L^{0}$-module $E$ and $G(n)$ be the group of permutations of $\{1,2,\cdots,n\}$. For any $\pi\in G(n)$, define
$y_{k}^{\pi}=\frac{1}{k}\sum_{i=1}^{k}x_{\pi(i)},\forall k\in\{1,2,\cdots,n\},$
$S_{\pi}=Conv_{L^{0}}(y_{1}^{\pi},y_{2}^{\pi},\cdots,y_{n}^{\pi}).$
Then, by applying the same method as in [11, Lemma 1.11], one can verify that $\{S_{\pi}:\pi\in G(n)\}$ is an $L^{0}$-simplicial subdivision of $S$, which is called the $L^{0}$-barycentric subdivision of $S$.
Example 3.2.
Let $S=Conv_{L^{0}}(\{x_{1},x_{2},x_{3}\})$ be a $2$-dimensional $L^{0}$-simplex in an $L^{0}$-module $E$. Let $y_{1,2}=\frac{1}{2}(x_{1}+x_{2})$, $y_{1,3}=\frac{1}{2}(x_{1}+x_{3})$ and $y_{2,3}=\frac{1}{2}(x_{2}+x_{3})$, it is easy to check that
$\mathscr{S}=\left\{\begin{array}[]{l}\begin{array}[]{l}Conv_{L^{0}}(\{x_{1},y_{1,2},y_{1,3}\}),Conv_{L^{0}}(\{x_{2},y_{1,2},y_{2,3}\}),\\ Conv_{L^{0}}(\{x_{3},y_{1,3},y_{2,3}\}),Conv_{L^{0}}(\{y_{1,2},y_{1,3},y_{2,3}\})\end{array}\end{array}\right\}$
is a family of $2$-dimensional $L^{0}$-simplexes.
(1). On the one hand, for any $x=\lambda_{1}x_{1}+\lambda_{2}x_{2}+\lambda_{3}x_{3}$ in $S$, where $\lambda_{1},\lambda_{2},\lambda_{3}\in L_{+}^{0}(\mathcal{F})$ and $\lambda_{1}+\lambda_{2}+\lambda_{3}=1$, let
$\begin{array}[]{ll}b_{1}=(\lambda_{1}\geq\frac{1}{2}),&\quad b_{2}=(\lambda_{2}\geq\frac{1}{2}),\\ b_{3}=(\lambda_{3}\geq\frac{1}{2}),&\quad b_{4}=(\lambda_{1}<\frac{1}{2})\wedge(\lambda_{2}<\frac{1}{2})\wedge(\lambda_{3}<\frac{1}{2})\end{array}$
and further let
$\begin{array}[]{ll}a_{1}=b_{1},&\quad a_{2}=b_{2}\wedge b_{1}^{c},\\ a_{3}=b_{3}\wedge(b_{1}\vee b_{2})^{c},&\quad a_{4}=b_{4},\end{array}$
then $\{a_{1},a_{2},a_{3},a_{4}\}\in p(1)$ and we have
$x$ $=\lambda_{1}x_{1}+\lambda_{2}x_{2}+\lambda_{3}x_{3}$
$=I_{a_{1}}[(\lambda_{1}-\lambda_{2}-\lambda_{3})x_{1}+2\lambda_{2}y_{1,2}+2\lambda_{3}y_{1,3}]+$
$\quad I_{a_{2}}[(\lambda_{2}-\lambda_{1}-\lambda_{3})x_{2}+2\lambda_{1}y_{1,2}+2\lambda_{3}y_{2,3}]+$
$\quad I_{a_{3}}[(\lambda_{3}-\lambda_{1}-\lambda_{2})x_{3}+2\lambda_{1}y_{1,3}+2\lambda_{2}y_{2,3}]+$
$\quad I_{a_{4}}[(\lambda_{1}+\lambda_{2}-\lambda_{3})y_{1,2}+(\lambda_{1}+\lambda_{3}-\lambda_{2})y_{1,3}+(\lambda_{2}+\lambda_{3}-\lambda_{1})y_{2,3}]$
$\in I_{a_{1}}Conv_{L^{0}}(\{x_{1},y_{1,2},y_{1,3}\})+I_{a_{2}}Conv_{L^{0}}(\{x_{2},y_{1,2},y_{2,3}\})+$
$\quad I_{a_{3}}Conv_{L^{0}}(\{x_{3},y_{1,3},y_{2,3}\})+I_{a_{4}}Conv_{L^{0}}(\{y_{1,2},y_{1,3},y_{2,3}\})$
$\subseteq\sigma\{\bigcup\mathscr{S}\},$
which implies that $S\subseteq\sigma\{\bigcup\mathscr{S}\}$. On the other hand, since $S$ is both $\sigma$-stable and $L^{0}$-convex, it follows that $\sigma\{\bigcup\mathscr{S}\}\subseteq S$. Therefore, $S=\sigma\{\bigcup\mathscr{S}\}$.
(2). For any $z\in Conv_{L^{0}}(\{x_{1},y_{1,2},y_{1,3}\})\cap Conv_{L^{0}}(\{y_{1,2},y_{1,3},y_{2,3}\})$, let
$z=\xi_{1}x_{1}+\xi_{2}y_{1,2}+\xi_{3}y_{1,3}=\eta_{1}y_{2,3}+\eta_{2}y_{1,3}+\eta_{3}y_{1,2},$
where $\{\xi_{1},\xi_{2},\xi_{3}\}\subseteq L_{+}^{0}(\mathcal{F})$, $\{\eta_{1},\eta_{2},\eta_{3}\}\subseteq L_{+}^{0}(\mathcal{F})$ and
$\xi_{1}+\xi_{2}+\xi_{3}=\eta_{1}+\eta_{2}+\eta_{3}=1,$
we have
$z=(\xi_{1}+\frac{\xi_{2}}{2}+\frac{\xi_{3}}{2})x_{1}+\frac{\xi_{2}}{2}x_{2}+\frac{\xi_{3}}{2}x_{3}=\frac{\eta_{2}+\eta_{3}}{2}x_{1}+\frac{\eta_{1}+\eta_{3}}{2}x_{2}+\frac{\eta_{1}+\eta_{2}}{2}x_{3}.$
Since $\{x_{1},x_{2},x_{3}\}$ is $L^{0}$-affinely independent,
[…]
then
[…]
which implies that $y\in Conv_{L^{0}}(\{y_{1,2},y_{1,3}\})$, and hence
$Conv_{L^{0}}(\{x_{1},y_{1,2},y_{1,3}\})\cap Conv_{L^{0}}(\{y_{1,2},y_{1,3},y_{2,3}\})=Conv_{L^{0}}(\{y_{1,2},y_{1,3}\}).$
Similarly, one can prove that
$Conv_{L^{0}}(\{x_{2},y_{1,2},y_{2,3}\})\cap Conv_{L^{0}}(\{y_{1,2},y_{1,3},y_{2,3}\})=Conv_{L^{0}}(\{y_{1,2},y_{2,3}\}),$
$Conv_{L^{0}}(\{x_{3},y_{1,3},y_{2,3}\})\cap Conv_{L^{0}}(\{y_{1,2},y_{1,3},y_{2,3}\})=Conv_{L^{0}}(\{y_{1,3},y_{2,3}\}),$
$Conv_{L^{0}}(\{x_{1},y_{1,2},y_{1,3}\})\cap Conv_{L^{0}}(\{x_{2},y_{1,2},y_{2,3}\})=\{y_{1,2}\},$
$Conv_{L^{0}}(\{x_{1},y_{1,2},y_{1,3}\})\cap Conv_{L^{0}}(\{x_{3},y_{1,3},y_{2,3}\})=\{y_{1,3}\}$
and
$Conv_{L^{0}}(\{x_{2},y_{1,2},y_{2,3}\})\cap Conv_{L^{0}}(\{x_{3},y_{1,3},y_{2,3}\})=\{y_{2,3}\}.$
Therefore, $\mathscr{S}$ satisfies (2) of Definition 2.6.
(3). Similar to the argument presented in (2), one can readily verify that
$\mathscr{S}^{\prime}=\left\{\begin{array}[]{l}\begin{array}[]{l}Conv(\{x_{1},y_{1,2},y_{1,3}\}),Conv(\{x_{2},y_{1,2},y_{2,3}\}),\\ Conv(\{x_{3},y_{1,3},y_{2,3}\}),Conv(\{y_{1,2},y_{1,3},y_{2,3}\})\end{array}\end{array}\right\}$
is a simplicial subdivision of $S^{\prime}$.
To sum up, $\mathscr{S}$ is an $L^{0}$-simplicial subdivision of $S$.
To prove Theorem 2.9, we need the following Lemma 3.3, Lemma 3.4 and Theorem 3.5. Although (2) of Lemma 3.3 is almost clear since each $S_{i}$ is $\sigma$-stable, we still give its proof for the reader’s convenience.
Lemma 3.3.
Let $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ be an $(n-1)$-dimensional $L^{0}$-simplex in an $L^{0}$-module $E$. Then we have the following statements:
  1. (1)
    For any sequence $\{S^{m}=Conv_{L^{0}}(\{x_{1}^{m},x_{2}^{m},\cdots,x_{n}^{m}\}),m\in\mathbb{N}\}$ of $(n-1)$-dimensional $L^{0}$-simplexes contained in $S$ and any $\{a_{m},m\in\mathbb{N}\}\in p(1)$, $\sum_{m=1}^{\infty}I_{a_{m}}S^{m}$ is an $(n-1)$-dimensional $L^{0}$-simplex contained in $S$ with vertexes $\sum_{m=1}^{\infty}I_{a_{m}}x_{1}^{m},\sum_{m=1}^{\infty}I_{a_{m}}x_{2}^{m},\cdots,$ $\sum_{m=1}^{\infty}I_{a_{m}}x_{n}^{m}$.
  2. (2)
    For any $L^{0}$-simplicial subdivision $\mathscr{S}=\{S_{i},i=1\sim I\}$ of $S$,
    $S=\bigcup\{\sum_{i=1}^{I}I_{b_{i}}S_{i}:\{b_{i},i=1\sim I\}\in p(1)\}.$
Proof.
(1). First, we prove that
$\sum_{m=1}^{\infty}I_{a_{m}}S^{m}=Conv_{L^{0}}(\{\sum_{m=1}^{\infty}I_{a_{m}}x_{1}^{m},\sum_{m=1}^{\infty}I_{a_{m}}x_{2}^{m},\cdots,\sum_{m=1}^{\infty}I_{a_{m}}x_{n}^{m}\}).$
For any $\{\lambda_{i},i=1\sim n\}\subseteq L_{+}^{0}(\mathcal{F})$ with $\sum_{i=1}^{n}\lambda_{i}=1$, we have
$\sum_{i=1}^{n}\lambda_{i}(\sum_{m=1}^{\infty}I_{a_{m}}x_{i}^{m})=\sum_{m=1}^{\infty}I_{a_{m}}(\sum_{i=1}^{n}\lambda_{i}x_{i}^{m})\in\sum_{m=1}^{\infty}I_{a_{m}}S^{m},$
which implies that
$Conv_{L^{0}}(\{\sum_{m=1}^{\infty}I_{a_{m}}x_{1}^{m},\sum_{m=1}^{\infty}I_{a_{m}}x_{2}^{m},\cdots,\sum_{m=1}^{\infty}I_{a_{m}}x_{n}^{m}\})\subseteq\sum_{m=1}^{\infty}I_{a_{m}}S^{m}.$
On the other hand, let $\sum_{i=1}^{n}\lambda_{i}^{m}x_{i}^{m}\in S^{m}$ for each $m\in\mathbb{N}$, where $\{\lambda_{i}^{m},i=1\sim n\}\subseteq L_{+}^{0}(\mathcal{F})$ and $\sum_{i=1}^{n}\lambda_{i}^{m}=1$, since $\{\sum_{m=1}^{\infty}I_{a_{m}}\lambda_{i}^{m},i=1\sim n\}\subseteq L_{+}^{0}(\mathcal{F})$ and
$\sum_{i=1}^{n}(\sum_{m=1}^{\infty}I_{a_{m}}\lambda_{i}^{m})=\sum_{m=1}^{\infty}I_{a_{m}}(\sum_{i=1}^{n}\lambda_{i}^{m})=1,$
we have
$\sum_{m=1}^{\infty}I_{a_{m}}(\sum_{i=1}^{n}\lambda_{i}^{m}x_{i}^{m})$ $=\sum_{i=1}^{n}(\sum_{m=1}^{\infty}I_{a_{m}}\lambda_{i}^{m}x_{i}^{m})$
$=\sum_{i=1}^{n}(\sum_{m=1}^{\infty}I_{a_{m}}\lambda_{i}^{m})(\sum_{m=1}^{\infty}I_{a_{m}}x_{i}^{m})$
$\in Conv_{L^{0}}(\{\sum_{m=1}^{\infty}I_{a_{m}}x_{1}^{m},\sum_{m=1}^{\infty}I_{a_{m}}x_{2}^{m},\cdots,\sum_{m=1}^{\infty}I_{a_{m}}x_{n}^{m}\}),$
which implies that
$\sum_{m=1}^{\infty}I_{a_{m}}S^{m}\subseteq Conv_{L^{0}}(\{\sum_{m=1}^{\infty}I_{a_{m}}x_{1}^{m},\sum_{m=1}^{\infty}I_{a_{m}}x_{2}^{m},\cdots,\sum_{m=1}^{\infty}I_{a_{m}}x_{n}^{m}\}).$
Second, we prove that $\{\sum_{m=1}^{\infty}I_{a_{m}}x_{1}^{m},\sum_{m=1}^{\infty}I_{a_{m}}x_{2}^{m},\cdots,\sum_{m=1}^{\infty}I_{a_{m}}x_{n}^{m}\}$ is $L^{0}$-affinely independent. Let $\{\mu_{i},i=1\sim n\}\subseteq L^{0}(\mathcal{F},\mathbb{K})$ with $\sum_{i=1}^{n}\mu_{i}=0$ and
$\sum_{i=1}^{n}\mu_{i}(\sum_{m=1}^{\infty}I_{a_{m}}x_{i}^{m})=0.$
For any $m\in\mathbb{N}$, since
$\sum_{i=1}^{n}(I_{a_{m}}\mu_{i})=I_{a_{m}}(\sum_{i=1}^{n}\mu_{i})=0,$
$\sum_{i=1}^{n}(I_{a_{m}}\mu_{i})x_{i}^{m}=I_{a_{m}}[\sum_{i=1}^{n}\mu_{i}(\sum_{m=1}^{\infty}I_{a_{m}}x_{i}^{m})]=0$
and $S^{m}$ is an $(n-1)$-dimensional $L^{0}$-simplex, then
$I_{a_{m}}\mu_{i}=0,\forall i\in\{1,2,\cdots,n\}.$
It follows from the arbitrariness of $m$ that
$\mu_{i}=0,\forall i\in\{1,2,\cdots,n\}.$
(2). Since $S$ is $\sigma$-stable, for any $\{b_{i},i=1\sim I\}\in p(1)$, we have
$\sum_{i=1}^{I}I_{b_{i}}S_{i}\subseteq\sum_{i=1}^{I}I_{b_{i}}(\bigcup_{i\in I}S_{i})\subseteq\sum_{i=1}^{I}I_{b_{i}}S=S,$
which implies that
$\bigcup\{\sum_{i=1}^{I}I_{b_{i}}S_{i}:\{b_{i},i=1\sim I\}\in p(1)\}\subseteq S.$
On the other hand, for any $x\in S$, since $S=\sigma(\bigcup_{i=1}^{I}S_{i})$, there exist $\{a_{n},n\in\mathbb{N}\}\in p(1)$ and a sequence $\{x_{n},n\in\mathbb{N}\}$ in $\bigcup_{i=1}^{I}S_{i}$ such that $x=\sum_{n=1}^{\infty}I_{a_{n}}x_{n}$. Let
$N_{i}=\{n\in\mathbb{N}:x_{n}\in S_{i}\},\forall i\in\{1,2,\cdots,I\}.$
For each $i\in\{1,2,\cdots,I\}$, since $S_{i}$ is $\sigma$-stable, arbitrarily fixing a $z_{i}\in S_{i}$, then there exists $y_{i}\in S_{i}$ such that $I_{a_{n}}x_{n}=I_{a_{n}}y_{i}$ for any $n\in N_{i}$ and $I_{(\bigvee_{n\in N_{i}}a_{n})^{c}}z_{i}=I_{(\bigvee_{n\in N_{i}}a_{n})^{c}}y_{i}$. Let
$b_{i}=\bigvee_{n\in N_{i}}a_{n},\forall i\in\{1,2,\cdots,I\},$
it is clear that $\{b_{i},i=1\sim I\}\in p(1)$ and
$x$ $=\sum_{n=1}^{\infty}I_{a_{n}}x_{n}$
$=\sum_{i=1}^{I}I_{b_{i}}y_{i}$
$\in\sum_{i=1}^{I}I_{b_{i}}S_{i}$
$\subseteq\bigcup\{\sum_{i=1}^{I}I_{b_{i}}S_{i}:\{b_{i},i=1\sim I\}\in p(1)\}.$
As a direct application of Lemma 5.4 of [26], Lemma 3.4 below gives an interesting connection between the set of $L^{0}$-extreme points of an $L^{0}$-simplicial subdivision and the set of extreme points of its counterpart simplicial subdivision.
Lemma 3.4.
Let $E$ and $S$ be the same as in Lemma 3.3, and $\mathscr{S}=\{S_{i},i=1\sim I\}$ be an $L^{0}$-simplicial subdivision of $S$. Then we have the following statements:
  1. (1)
    $ext_{L^{0}}(S)=\sigma(ext(S^{\prime}))$.
  2. (2)
    $ext_{L^{0}}(\mathscr{S})=\sigma(ext(\mathscr{S}^{\prime}))$.
Proof.
(1). By Lemma 5.4 of [26], we have
$ext_{L^{0}}(S)=\sigma(\{x_{1},x_{2},\cdots,x_{n}\}),$
which, combined with the fact $ext(S^{\prime})=\{x_{1},x_{2},\cdots,x_{n}\}$, implies that
$ext_{L^{0}}(S)=\sigma(ext(S^{\prime})).$
(2). Since $ext_{L^{0}}(S_{i})=\sigma(ext(S^{\prime}_{i}))$ for each $i\in\{1,2,\cdots,I\}$, we have
$ext_{L^{0}}(\mathscr{S})$ $=\sigma(\bigcup_{i=1}^{I}ext_{L^{0}}(S_{i}))$
$=\sigma(\bigcup_{i=1}^{I}\sigma(ext(S^{\prime}_{i})))$
$=\sigma(\bigcup_{i=1}^{I}ext(S^{\prime}_{i}))$
$=\sigma(ext(\mathscr{S}^{\prime})).$
Throughout this paper, for any $\xi=\sum_{i=1}^{n}I_{a_{i}}i\in L^{0}(\mathcal{F},\{1,2,\cdots,n\})$, in such an expression of $\xi$, from now on we always assume that $a_{i}=(\xi=i)$ for each $i\in\{1,2,\cdots,n\}$ (then it is clear that $\{a_{i},i=1\sim n\}\in p(1)$), we will also always use $\mathcal{N}_{\xi}$ for the set $\{i:a_{i}>0\}$.
For any $\{a_{n},n\in\mathbb{N}\},~{}\{b_{m},m\in\mathbb{N}\}\in p(1)$, it is easy to check that $\{a_{n}\wedge b_{m},n,m\in\mathbb{N}\}\in p(1)$, which is referred to as the refinement of $\{a_{n},n\in\mathbb{N}\}$ and $\{b_{m},m\in\mathbb{N}\}$. The refinement of any finite collection of partitions can be defined inductively.
Theorem 3.5 below establishes a key connection between the set of proper $L^{0}$-labeling functions of $\mathscr{S}$ and the set of proper labeling functions of $\mathscr{S^{\prime}}$. In particular, (3) of Theorem 3.5 below gives the representation theorem of a proper $L^{0}$-labeling function by usual proper labeling functions, which considerably simplifies the proof of Theorem 2.9 below.
Theorem 3.5.
Let $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ be an $(n-1)$-dimensional $L^{0}$-simplex in an $L^{0}$-module $E$ and $\mathscr{S}$ be an $L^{0}$-simplicial subdivision of $S$. Then we have the following statements:
  1. (1)
    For any proper $L^{0}$-labeling function $\phi$ of $\mathscr{S}$,
    $\mathcal{N}_{\phi(y)}\subseteq\chi(y),\forall y\in ext(\mathscr{S}^{\prime}).$
  2. (2)
    For any proper $L^{0}$-labeling function $\phi$ of $\mathscr{S}$ and any $y\in ext(\mathscr{S}^{\prime})$, there exists a family $\{\psi_{i}:i\in\mathcal{N}_{\phi(y)}\}$ of proper labeling functions of $\mathscr{S^{\prime}}$ with $\psi_{i}(y)=i$ for each $i\in\mathcal{N}_{\phi(y)}$, namely
    $\phi(y)=\sum_{i\in\mathcal{N}_{\phi(y)}}I_{(\phi(y)=i)}\psi_{i}(y).$
  3. (3)
    Denote $ext(\mathscr{S}^{\prime})=\{y_{k}:k=1\sim K\}$ and let $\phi$ be a proper $L^{0}$-labeling function of $\mathscr{S}$. Then there exist a finite family $\{\psi_{m},m=1\sim M\}$ of proper labeling functions of $\mathscr{S}^{\prime}$ and $\{a_{m},m=1\sim M\}\in p(1)$ such that
    $\phi(y_{k})=\sum_{m=1}^{M}I_{a_{m}}\psi_{m}(y_{k})~{}\text{for~{}any}~{}k\in\{1,2,\cdots,K\}.$
    More generally, since $\phi$ is $\sigma$-stable, we also have
    $\phi(y)=\sum_{k=1}^{K}I_{b_{k}}\sum_{m=1}^{M}I_{a_{m}}\psi_{m}(y_{k})$
    for any $y=\sum_{k=1}^{K}I_{b_{k}}y_{k}\in ext_{L^{0}}(\mathscr{S})$ with $\{b_{k},k=1\sim K\}\in p(1)$.
Proof.
(1). For any given $y\in ext(\mathscr{S}^{\prime})\subseteq S^{\prime}$, there exists $\{\alpha_{i},i=1\sim n\}\subseteq[0,1]$ with $\sum_{i=1}^{n}\alpha_{i}=1$ such that
$y=\sum_{i=1}^{n}\alpha_{i}x_{i}\in S.$
For any $i\in\mathcal{N}_{\phi(y)}$, if $\alpha_{i}=0$, then $(\alpha_{i}=0)=1$, which, combined with
$(\phi(y)=i)\wedge(\alpha_{i}=0)=0,$
implies that $(\phi(y)=i)=0$. This contradicts the fact that $i\in\mathcal{N}_{\phi(y)}$. Hence, $\alpha_{i}>0$ and this implies that $i\in\chi(y)$. Thus,
$\mathcal{N}_{\phi(y)}\subseteq\chi(y).$
(2). For each $i\in\mathcal{N}_{\phi(y)}$, since $\mathcal{N}_{\phi(y)}\subseteq\chi(y)$, it is clear that there exists a proper labeling function $\psi_{i}$ of $\mathscr{S^{\prime}}$ such that $\psi_{i}(y)=i$, namely
$\phi(y)=\sum_{i\in\mathcal{N}_{\phi(y)}}I_{(\phi(y)=i)}\psi_{i}(y).$
(3). For each $k\in\{1,2,\cdots,K\}$, let $y_{k}=\sum_{i=1}^{n}\lambda_{i}^{k}x_{i}$, where $\{\lambda_{i}^{k},i=1\sim n\}\subseteq[0,1]$ are such that $\sum_{i=1}^{n}\lambda_{i}^{k}=1$. Since $\phi$ maps $ext_{L^{0}}(\mathscr{S})$ into $L^{0}(\mathcal{F},\{1,2,\cdots,n\})$, for each $k\in\{1,2,\cdots,K\}$, there exists $\{a_{i}^{k},i=1\sim n\}\in p(1)$ such that
$\phi(y_{k})=\sum_{i=1}^{n}I_{a_{i}^{k}}i.$
Now, we consider the refinement
$\{a_{i_{1}}^{1}\wedge a_{i_{2}}^{2}\wedge\cdots\wedge a_{i_{K}}^{K}:(i_{1},i_{2},\cdots,i_{K})\in\{1,2,\cdots,n\}^{K}\}$
of $\{a_{i}^{1},i=1\sim n\},\cdots,\{a_{i}^{K-1},i=1\sim n\}$ and $\{a_{i}^{K},i=1\sim n\}$, where $\{1,2,\cdots,n\}^{K}$ is the $K$-th Cartesian power of $\{1,2,\cdots,n\}$. Further, let
$I=\{(i_{1},i_{2},\cdots,i_{K})\in\{1,2,\cdots,n\}^{K}:a_{i_{1}}^{1}\wedge a_{i_{2}}^{2}\wedge\cdots\wedge a_{i_{K}}^{K}>0\},$
then
$\{a_{i_{1}}^{1}\wedge a_{i_{2}}^{2}\wedge\cdots\wedge a_{i_{K}}^{K}:(i_{1},i_{2},\cdots,i_{K})\in I\}\in p(1)$
and
$\mathcal{N}_{\phi(y_{k})}=\{P_{k}(i_{1},i_{2},\cdots,i_{K}):(i_{1},i_{2},\cdots,i_{K})\in I\},\forall k\in\{1,2,\cdots,K\},$
where $P_{k}(\cdot)$ denotes the projection of $\{1,2,\cdots,n\}^{K}$ onto the $k$-th coordinate.
For each $k\in\{1,2,\cdots,K\}$, let
$\mathcal{N}_{\phi(y_{k})}(\text{namely}~{}P_{k}(I))=\{p_{k,l}:1\leq l\leq M_{k}\},$
where $M_{k}$ denotes the cardinality of $P_{k}(I)$. Further, for each $l\in\{1,2,\cdots,M_{k}\}$, let
$I_{k,l}=\{(i_{1},i_{2},\cdots,i_{K})\in I:P_{k}(i_{1},i_{2},\cdots,i_{K})=p_{k,l}\},$
that is to say, $I_{k,l}=P_{k}^{-1}(p_{k,l})$, then it is clear that $I=\bigcup_{l=1}^{M_{k}}I_{k,l}$ and $\{I_{k,l}:1\leq l\leq M_{k}\}$ is a pairwise disjoint family. Now, by (2) as has been proved, there exists a family $\{\varphi_{k,l}:1\leq l\leq M_{k}\}$ of proper labeling functions of $\mathscr{S}^{\prime}$ such that $\varphi_{k,l}(y_{k})=p_{k,l}$ for each $l\in\{1,2,\cdots M_{k}\}$, namely
$\phi(y_{k})=\sum_{l=1}^{M_{k}}I_{(\phi(y_{k})=p_{k,l})}p_{k,l}=\sum_{l=1}^{M_{k}}I_{(\phi(y_{k})=p_{k,l})}\varphi_{k,l}(y_{k}).$
For each $k\in\{1,2,\cdots,K\}$ and each $l\in\{1,2,\cdots,M_{k}\}$, we define a proper labeling function $\psi_{(i_{1},i_{2},\cdots,i_{K})}^{k}$ of $\mathscr{S}^{\prime}$ for each $(i_{1},i_{2},\cdots,i_{K})\in I_{k,l}$ by $\psi_{(i_{1},i_{2},\cdots,i_{K})}^{k}=\varphi_{k,l}$, then one can easily see that
$\psi_{(i_{1},i_{2},\cdots,i_{K})}^{k}(y_{k})=\varphi_{k,l}(y_{k})=p_{k,l}.$
To sum up, since $I=\bigcup_{l=1}^{M_{k}}I_{k,l}$, we can eventually obtain a family
$\{\psi_{(i_{1},i_{2},\cdots,i_{K})}^{k}:k\in\{1,2,\cdots,K\}~{}\text{and}~{}(i_{1},i_{2},\cdots,i_{K})\in I\}$
of proper labeling functions of $\mathscr{S}^{\prime}$ such that each $\psi_{(i_{1},i_{2},\cdots,i_{K})}^{k}$ in this family satisfies that
$\psi^{k}_{(i_{1},i_{2},\cdots,i_{K})}(y_{k})=P_{k}(i_{1},i_{2},\cdots,i_{K}).$
Since, for each $k\in\{1,2,\cdots,K\}$ and each $l\in\{1,2,\cdots,M_{k}\}$,
$(\phi(y_{k})=p_{k,l})=\bigvee_{(i_{1},i_{2},\cdots,i_{K})\in I_{k,l}}a_{i_{1}}^{1}\wedge a_{i_{2}}^{2}\wedge\cdots\wedge a_{i_{K}}^{K},$
then
$I_{(\phi(y_{k})=p_{k,l})}=\sum_{(i_{1},i_{2},\cdots,i_{K})\in I_{k,l}}I_{a_{i_{1}}^{1}\wedge a_{i_{2}}^{2}\wedge\cdots\wedge a_{i_{K}}^{K}},$
and hence
$\phi(y_{k})$ $=\sum_{l=1}^{M_{k}}I_{(\phi(y_{k})=p_{k,l})}p_{k,l}$
$=\sum_{l=1}^{M_{k}}\sum_{(i_{1},i_{2},\cdots,i_{K})\in I_{k,l}}I_{a_{i_{1}}^{1}\wedge a_{i_{2}}^{2}\wedge\cdots\wedge a_{i_{K}}^{K}}p_{k,l}$
$=\sum_{l=1}^{M_{k}}\sum_{(i_{1},i_{2},\cdots,i_{K})\in I_{k,l}}I_{a_{i_{1}}^{1}\wedge a_{i_{2}}^{2}\wedge\cdots\wedge a_{i_{K}}^{K}}\psi_{(i_{1},i_{2},\cdots,i_{K})}^{k}(y_{k})$
$=\sum_{(i_{1},i_{2},\cdots,i_{K})\in I}I_{a_{i_{1}}^{1}\wedge a_{i_{2}}^{2}\wedge\cdots\wedge a_{i_{K}}^{K}}\psi_{(i_{1},i_{2},\cdots,i_{K})}^{k}(y_{k}).$
Now, for each $(i_{1},i_{2},\cdots,i_{K})\in I$, we define a mapping $\psi_{(i_{1},i_{2},\cdots,i_{K})}:ext(\mathscr{S^{\prime}})\rightarrow\{1,2,\cdots,n\}$ by
$\psi_{(i_{1},i_{2},\cdots,i_{K})}(y_{k})=\psi^{k}_{(i_{1},i_{2},\cdots,i_{K})}(y_{k}),\forall k\in\{1,2,\cdots,K\}.$
Then $\psi_{(i_{1},i_{2},\cdots,i_{K})}$ is a proper labeling function of $\mathscr{S}^{\prime}$. It follows that
$\phi(y_{k})=\sum_{(i_{1},i_{2},\cdots,i_{K})\in I}I_{a_{i_{1}}^{1}\wedge a_{i_{2}}^{2}\wedge\cdots\wedge a_{i_{K}}^{K}}\psi_{(i_{1},i_{2},\cdots,i_{K})}(y_{k}),\forall k\in\{1,2,\cdots,K\}.$
Furthermore, for any $y=\sum_{k=1}^{K}I_{b_{k}}y_{k}\in ext_{L^{0}}(\mathscr{S})$ with $\{b_{k},k=1\sim K\}\in p(1)$, we have
$\phi(\sum_{k=1}^{K}I_{b_{k}}y_{k})$ $=\sum_{k=1}^{K}I_{b_{k}}\phi(y_{k})$
$=\sum_{k=1}^{K}I_{b_{k}}(\sum_{(i_{1},i_{2},\cdots,i_{K})\in I}I_{a_{i_{1}}^{1}\wedge a_{i_{2}}^{2}\wedge\cdots\wedge a_{i_{K}}^{K}}\psi_{(i_{1},i_{2},\cdots,i_{K})}(y_{k})).$
Lemma 3.3, Lemma 3.4 and Theorem 3.5 show that $L^{0}$-simplexes, $L^{0}$-simplicial subdivisions and proper $L^{0}$-labeling functions are closely connected to their classical counterparts. These results motivate us to make full use of Lemma 2.8 to complete the proof of Theorem 2.9.
We can now prove Theorem 2.9.
Proof of Theorem 2.9.
Let $ext(\mathscr{S}^{\prime})=\{y_{k}:k=1\sim K\}$ and $y_{k}=\sum_{i=1}^{n}\lambda_{i}^{k}x_{i}$ with $\{\lambda_{i}^{k},i=1\sim n\}\subseteq[0,1]$ and $\sum_{i=1}^{n}\lambda_{i}^{k}=1$ for any $k\in\{1,2,\cdots,K\}$. By (3) of Theorem 3.5, there exist a finite family $\{\psi_{m},m=1\sim M\}$ of proper labeling functions of $\mathscr{S}^{\prime}$ and $\{a_{m},m=1\sim M\}\in p(1)$ such that
$\phi(y_{k})=\sum_{m=1}^{M}I_{a_{m}}\psi_{m}(y_{k}),~{}\forall k\in\{1,2,\dots,K\}.$
For each $\psi_{m}$, by Lemma 2.8, there exists $S^{\prime}_{m}\in\mathscr{S}^{\prime}$ such that $S^{\prime}_{m}$ is completely labeled by $\psi_{m}$. Let $S^{\prime}_{m}=Conv(\{y_{1}^{m},\cdots,y_{n}^{m}\})$, where $y_{j}^{m}\in ext(\mathscr{S}^{\prime})$ and
$\psi_{m}(y_{j}^{m})=j,\forall j\in\{1,2,\cdots,n\}.$ (3.1)
Further let $S_{m}=Conv_{L^{0}}(\{y_{1}^{m},\cdots,y_{n}^{m}\})$ for each $m\in\{1,2,\cdots,M\}$, then by (1) of Lemma 3.3,
$\sum_{m=1}^{M}I_{a_{m}}S_{m}=Conv_{L^{0}}(\{\sum_{m=1}^{M}I_{a_{m}}y_{1}^{m},\sum_{m=1}^{M}I_{a_{m}}y_{2}^{m},\cdots,\sum_{m=1}^{M}I_{a_{m}}y_{n}^{m}\})$
is an $(n-1)$-dimensional $L^{0}$-simplex contained in $S$. Furthermore, by (3.1), we have
$\phi(\sum_{m=1}^{M}I_{a_{m}}y_{j}^{m})$ $=\sum_{m=1}^{M}I_{a_{m}}\phi(y_{j}^{m})$
$=\sum_{m=1}^{M}I_{a_{m}}(\sum_{l=1}^{M}I_{a_{l}}\psi_{l}(y_{j}^{m}))$
$=\sum_{m=1}^{M}I_{a_{m}}\psi_{m}(y_{j}^{m})$
$=\sum_{m=1}^{M}I_{a_{m}}j$
$=j$
for any $j\in\{1,2,\cdots,n\}$. This implies that $\sum_{m=1}^{M}I_{a_{m}}S_{m}$ is completely labeled by $\phi$. ∎
4 Proof of Theorem 2.1
Although Theorem 2.1 is stated in $L^{0}(\mathcal{F},\mathbb{R}^{d})$, several key concepts that will be used in its proof originate from the theory of random normed modules. For clarity and convenience, we begin this section by recalling some basic notions from the theory of random normed modules, and we also restate several concepts previously introduced in Section 2 within the framework of random normed modules.
Definition 4.1 here is adopted from [19, 20] by following the traditional nomenclature of random metric spaces and random normed spaces (see [47, Chapters 9 and 15]).
Definition 4.1 ([19, 20]).
[…]
  1. (1)
    $\|\xi x\|=|\xi|\|x\|$ for any $\xi\in L^{0}(\mathcal{F},\mathbb{K})$ and any $x\in E$, where $\xi x$ stands for the module multiplication of $x$ by $\xi$ and $|\xi|$ is the equivalence class of $|\xi^{0}|$ for an arbitrarily chosen representative $\xi^{0}$ of $\xi$ (of course, $|\xi^{0}|$ is the function defined by $|\xi^{0}|(\omega)=|\xi^{0}(\omega)|$ for any $\omega\in\Omega$);
  2. (2)
    $\|x+y\|\leq\|x\|+\|y\|$ for any $x$ and $y$ in $E$;
  3. (3)
    $\|x\|=0$ implies $x=\theta$ (the null element of $E$).
[…]
It should be mentioned that the notion of an $L^{0}$-normed $L^{0}$-module, which is equivalent to that of an $RN$ module, was independently introduced by Gigli in [16] for the study of nonsmooth differential geometry on metric measure spaces, where the $L^{0}$-norm was called the pointwise norm.
Clearly, when $(\Omega,\mathcal{F},P)$ is trivial, namely =$\mathcal{F}=\{\Omega,\emptyset\}$, then an $RN$ module with base $(\Omega,\mathcal{F},P)$ just reduces to an ordinary normed space.
Let $(E,\|\cdot\|)$ be an $RN$ module $(E,\|\cdot\|)$ over $\mathbb{K}$ with base $(\Omega,\mathcal{F},P)$. For any real numbers $\varepsilon>0$ and $10<\lambda<$ 1, let
$U_{\theta}(\varepsilon,\lambda)=\{x\in E:P\{\omega\in\Omega:\|x\|(\omega)<\varepsilon\}>1-\lambda\}.$
[…]
$B(x,r)=\{y\in E:\|y\|<r~{}\text{on}~{}\Omega\},$
[…]
Example 4.2.
Let $L^{0}(\mathcal{F},\mathbb{K}^{d})$ be the linear space of equivalence classes of random vectors from $(\Omega,\mathcal{F},P)$ to $\mathbb{K}^{d}$, then $L^{0}(\mathcal{F},\mathbb{K}^{d})$ forms an $L^{0}$-module under the module multiplication $\cdot:L^{0}(\mathcal{F},\mathbb{K})\times L^{0}(\mathcal{F},\mathbb{K}^{d})$ defined by
$\lambda\cdot x=(\lambda\xi_{1},\lambda\xi_{2},\cdots,\lambda\xi_{d}),\forall\lambda\in L^{0}(\mathcal{F},\mathbb{K}),\forall x=(\xi_{1},\xi_{2},\cdots,\xi_{d})\in L^{0}(\mathcal{F},\mathbb{K}^{d}).$
Define the $L^{0}$-inner product $\langle\cdot,\cdot\rangle:L^{0}(\mathcal{F},\mathbb{K}^{d})\times L^{0}(\mathcal{F},\mathbb{K}^{d})$ by
$\langle x,y\rangle=\sum_{i=1}^{d}x_{i}\bar{y}_{i},\forall x=(x_{1},x_{2},\cdots,x_{d}),y=(y_{1},y_{2},\cdots,y_{d})\in L^{0}(\mathcal{F},\mathbb{K}^{d}),$
where $\bar{y}_{i}$ denotes the complex conjugate of $y_{i}$ for each $i\in\{1,2,\cdots,n\}$. Then $(L^{0}(\mathcal{F},\mathbb{K}^{d}),\langle\cdot,\cdot\rangle)$ is a $\mathcal{T}_{\varepsilon,\lambda}$-complete random inner product module (briefly, a $\mathcal{T}_{\varepsilon,\lambda}$-complete $RIP$ module, see [22] for details). let $\|x\|=\sqrt{\langle x,x\rangle}$ for any $x\in L^{0}(\mathcal{F},\mathbb{K}^{d})$, then $(L^{0}(\mathcal{F},\mathbb{K}^{d}),\|\cdot\|)$ is a $\mathcal{T}_{\varepsilon,\lambda}$-complete $RN$ module.
In fact, $L^{0}(\mathcal{F},\mathbb{K}^{d})$ is also $\mathcal{T}_{c}$-complete. Generally, according to Theorem 3.18 of [23], an $RN$ module is $\mathcal{T}_{\varepsilon,\lambda}$-complete iff it is both $\mathcal{T}_{c}$-complete and $\sigma$-stable.
The randomized version of the Bolzano-Weierstrass theorem [31] states that for any a.s. bounded sequence $\{\xi^{0}_{n},n\in\mathbb{N}\}$ of random vectors from $(\Omega,\mathcal{F},P)$ to $\mathbb{R}^{d}$ (namely $\sup\{|\xi^{0}_{n}(\omega)|:n\in\mathbb{N}\}<+\infty$ $a.s.$) there exists a sequence $\{n_{k}^{0},k\in\mathbb{N}\}$ of positive integer-valued random variables on $(\Omega,\mathcal{F},P)$ such that $\{n_{k}^{0}(\omega),k\in\mathbb{N}\}$ is strictly increasing for each $\omega\in\Omega$ and $\{\xi^{0}_{n_{k}^{0}},k\in\mathbb{N}\}$ converges a.s., where $\xi^{0}_{n_{k}^{0}}(\omega)=\xi^{0}_{n_{k}^{0}(\omega)}(\omega)=\sum^{\infty}_{l=1}I_{(n_{k}^{0}=l)}(\omega)\xi^{0}_{l}(\omega)$ for any $k\in\mathbb{N}$ and any $\omega\in\Omega$ (as usual, $(n_{k}^{0}=l)$ denotes the set $\{\omega\in\Omega:n_{k}^{0}(\omega)=l\}$ for any $k$ and $l$ in $\mathbb{N}$). Motivated by this fundamental result, Guo et al. introduced the following notion of random sequential compactness in $RN$ modules.
Definition 4.3 ([26]).
Let $(E,\|\cdot\|)$ be an $RN$ module over $\mathbb{K}$ with base $(\Omega,\mathcal{F},P)$ and $G$ a nonempty subset such that $G$ is contained in a $\sigma$-stable subset $H$ of $E$. Given a sequence $\{x_{n},n\in\mathbb{N}\}$ in $G$, a sequence $\{y_{k},k\in\mathbb{N}\}$ in $H$ is called a random subsequence of $\{x_{n},n\in\mathbb{N}\}$ if there exists a sequence $\{n_{k},k\in\mathbb{N}\}$ in $L^{0}(\mathcal{F},\mathbb{N})$ such that the following two conditions are satisfied:
  1. (1)
    $n_{k}<n_{k+1}$ on $\Omega$;
  2. (2)
    $y_{k}=x_{n_{k}}:=\sum^{\infty}_{l=1}I_{(n_{k}=l)}x_{l}$ for each $k\in\mathbb{N}$.
Further, $G$ is said to be random (resp., relatively) sequentially compact if there exists a random subsequence $\{y_{k},k\in\mathbb{N}\}$ of $\{x_{n},n\in\mathbb{N}\}$ for any sequence $\{x_{n},n\in\mathbb{N}\}$ in $G$ such that $\{y_{k},k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to some element in $G$ (resp., in $E$).
Recall that a set $G$ of an $RN$ module with base $(\Omega,\mathcal{F},P)$ is said to be a.s. bounded if $\bigvee_{x\in G}\|x\|\in L_{+}^{0}(\mathcal{F})$.
Remark 4.4.
(1) In Definition 4.3, we require $n_{k}$ to be an element in $L^{0}(\mathcal{F},\mathbb{N})$, rather than a positive integer-valued random variable as in [26, Definition 2.10]. This choice is made for simplifying written expressions, and it is easy to check that the two formulations are essentially equivalent. (2) In the sense of Definition 4.3, the randomized Bolzano-Weierstrass theorem can be restated as: any a.s. bounded nonempty subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is random relatively sequentially compact. Moreover, as pointed out in [26], if $G$ is $\sigma$-stable, then $G$ is random relatively sequentially compact iff $\bar{G}_{\varepsilon,\lambda}$ (the closure of $G$ under $\mathcal{T}_{\varepsilon,\lambda}$) is random sequentially compact. Consequently, any a.s. bounded $\sigma$-stable nonempty $\mathcal{T}_{\varepsilon,\lambda}$-closed subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ (for example, an $L^{0}$-simplex) must be random sequentially compact. (3) For any $(n-1)$-dimensional $L^{0}$-simplex $S$ of an $RN$ module, as pointed out by [26, Lemma 5.5], there exists an $L^{0}$-affine $\mathcal{T}_{\varepsilon,\lambda}$-isomorphism $T$ from $\Delta_{n-1}$ to $S$, where $\Delta_{n-1}:=Conv_{L^{0}}(\{e_{1},e_{2},\cdots,e_{n}\})$ is the standard $(n-1)$-dimensional $L^{0}$-simplex of $L^{0}(\mathcal{F},\mathbb{R}^{n})$. Since $\Delta_{n-1}$ is random sequentially compact as discussed above, and $T$ is $\sigma$-stable and $\mathcal{T}_{\varepsilon,\lambda}$-continuous, then $S$ is random sequentially compact.
Definition 4.5 ([26]).
Let $(E_{1},\|\cdot\|_{1})$ and $(E_{2},\|\cdot\|_{2})$ be two $RN$ modules over the same scalar field $\mathbb{K}$ with the same base $(\Omega,\mathcal{F},P)$, and $G_{1}$ and $G_{2}$ be two nonempty subsets of $E_{1}$ and $E_{2}$, respectively. A mapping $T:G_{1}\rightarrow G_{2}$ is said to be:
  1. (1)
    $\mathcal{T}_{\varepsilon,\lambda}$-continuous if $T$ is a continuous mapping from $(G_{1},\mathcal{T}_{\varepsilon,\lambda})$ to $(G_{2},\mathcal{T}_{\varepsilon,\lambda})$.
  2. (2)
    $\mathcal{T}_{c}$-continuous if $T$ is a continuous mapping from $(G_{1},\mathcal{T}_{c})$ to $(G_{2},\mathcal{T}_{c})$.
  3. (3)
    $a.s.$ sequentially continuous at $x_{0}\in G_{1}$ if $\{\|T(x_{n})-T(x_{0})\|_{2},n\in\mathbb{N}\}$ converges $a.s.$ to 00 for any sequence $\{x_{n},n\in\mathbb{N}\}$ in $G_{1}$ such that $\{\|x_{n}-x_{0}\|,n\in\mathbb{N}\}$ converges $a.s.$ to 00. Further, $T$ is said to be $a.s.$ sequentially continuous if $T$ is $a.s.$ sequentially continuous at any point in $G_{1}$.
  4. (4)
    random sequentially continuous at $x_{0}\in G_{1}$ if $G_{1}$ is $\sigma$-stable and for any sequence $\{x_{n},n\in\mathbb{N}\}$ in $G_{1}$ convergent in $\mathcal{T}_{\varepsilon,\lambda}$ to $x_{0}$ there exists a random subsequence $\{x_{n_{k}},k\in\mathbb{N}\}$ of $\{x_{n},n\in\mathbb{N}\}$ such that $\{T(x_{n_{k}}),k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $T(x_{0})$. Further, $T$ is said to be random sequentially continuous if $T$ is random sequentially continuous at any point in $G_{1}$.
Remark 4.6.
It is clear that a.s. sequential continuity implies $\mathcal{T}_{\varepsilon,\lambda}$-continuity. Moreover, when $G_{1}$ is $\sigma$-stable, the $\mathcal{T}_{\varepsilon,\lambda}$-continuity of a mapping $T$ from $G_{1}$ to $G_{2}$ necessarily implies random sequential continuity of $T$. Furthermore, [26, Lemma 4.3] shows that a $\sigma$-stable mapping $T$ of $G_{1}$ to $G_{2}$ is random sequentially continuous iff it is $\mathcal{T}_{c}$-continuous.
For any $(n-1)$-dimensional $L^{0}$-simplex $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ in an $RN$ module $E$ and any $L^{0}$-simplicial subdivision $\mathscr{S}=\{S_{i}:i\in I\}$ of $S$, the random diameter of $\mathscr{S}$ is defined by
$diam(\mathscr{S})=\bigvee_{i\in I}diam(S_{i}),$
where $diam(S_{i}):=\bigvee\{\|x-y\|:x,y\in S_{i}\}$ is called the random diameter of $S_{i}$ for each $i\in I$.
Lemma 4.7.
Let $(E,\|\cdot\|)$ be an $RN$ module over $\mathbb{K}$ with base $(\Omega,\mathcal{F},P)$, $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ an $(n-1)$-dimensional $L^{0}$-simplex in $E$ and $\mathscr{S}$ an $L^{0}$-barycentric subdivision of $S$. Then
$diam(\mathscr{S})\leq\frac{n-1}{n}diam(S).$
Proof.
First, we claim that $diam(G)=\bigvee_{k,l=1}^{m}\|y_{i}-y_{j}\|$ for any $(n-1)$-dimensional $L^{0}$-simplex $G=Conv_{L^{0}}(\{y_{1},y_{2},\cdots,y_{n}\})$ in $E$. In fact, it is clear that $\bigvee_{k,l=1}^{m}\|y_{k}-y_{l}\|\leq diam(G)$. On the other hand, for any $x=\sum_{i=1}^{m}\lambda_{i}y_{i}$ and $y=\sum_{i=1}^{m}\mu_{i}y_{i}$ in $G$, where $\lambda_{i},\mu_{i}\in L_{+}^{0}(\mathcal{F})$ for any $i\in\{1,2,\cdots,m\}$ and $\sum_{i=1}^{m}\lambda_{i}=\sum_{i=1}^{m}\mu_{i}=1$, we have
$\|x-y\|$ $=\|\sum_{i=1}^{m}\lambda_{i}y_{i}-\sum_{i=1}^{m}\mu_{i}y_{i}\|$
$=\|\sum_{i,j=1}^{m}\lambda_{i}\mu_{j}(y_{i}-y_{j})\|$
$\leq\sum_{i,j=1}^{m}\lambda_{i}\mu_{j}(\bigvee_{k,l=1}^{m}\|y_{k}-y_{l}\|)$
$=\bigvee_{k,l=1}^{m}\|y_{k}-y_{l}\|.$
Second, for any $\pi\in G(n)$ and any $k,l\in\{1,2,\cdots,n\}$, without loss of generality, we can assume that $k>l$, then
$\|y_{k}^{\pi}-y_{l}^{\pi}\|$ $=\|\frac{1}{k}\sum_{i=1}^{k}x_{\pi(i)}-\frac{1}{l}\sum_{j=1}^{l}x_{\pi(j)}\|$
$=\frac{k-l}{k}\|\frac{1}{l}\sum_{i=1}^{l}x_{\pi(i)}-\frac{1}{k-l}\sum_{j=l+1}^{k}x_{\pi(j)}\|$
$\leq\frac{n-1}{n}\|\frac{1}{l}\sum_{i=1}^{l}x_{\pi(i)}-\frac{1}{k-l}\sum_{j=l+1}^{k}x_{\pi(j)}\|$
$\leq\frac{n-1}{n}diam(S),$
which implies that
$diam(S_{\pi})=\bigvee_{k,l=1}^{n}\|y_{k}^{\pi}-y_{l}^{\pi}\|\leq\frac{n-1}{n}diam(S).$
It follows that
$diam(\mathscr{S})=\bigvee_{\pi\in G(n)}diam(S_{\pi})\leq\frac{n-1}{n}diam(S).$
The key idea of the proof of classical Brouwer fixed point theorem for a continuous self-mapping on a simplex $S^{\prime}$ lies in constructing a proper labeling function $\psi$ on a given simplicial subdivision $\mathscr{S}^{\prime}$ of $S^{\prime}$ that is compatible with the continuous self-mapping on $S^{\prime}$, see Theorem 6.1 of [1] for details. Lemma 4.8 below provides a method for constructing such an $L^{0}$-labeling function in connection with an $(n-1)$-dimensional $L^{0}$-simplex $S$ in an $RN$ module, which improves the formulation of Lemma 2.2 of [11] in that the single proper $L^{0}$-labeling function provided here can be applied to the $L^{0}$-simplicial subdivision of every $(n-1)$-dimensional $L^{0}$-simplex contained in $S$.
Lemma 4.8.
Let $(E,\|\cdot\|)$ be an $RN$ module with base $(\Omega,\mathcal{F},P)$, $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ an $(n-1)$-dimensional $L^{0}$-simplex in $E$ and $f:S\rightarrow S$ be a $\sigma$-stable mapping. Then there exists a mapping $\phi:S\rightarrow L^{0}(\mathcal{F},\{1,2,\cdots,n\})$ such that the following conditions are satisfied:
  1. (1)
    For any $x=\sum_{i=1}^{n}\lambda_{i}^{x}x_{i}\in S$, further let $f(x)=\sum_{i=1}^{n}\mu_{i}^{x}x_{i}$, where $\lambda_{i}^{x},\mu_{i}^{x}\in L_{+}^{0}(\mathcal{F})$ for any $i\in\{1,2,\cdots,n\}$ and $\sum_{i=1}^{n}\lambda_{i}^{x}=\sum_{i=1}^{n}\mu_{i}^{x}=1$, it holds that
    $(\phi(x)=i)\wedge[(\lambda_{i}^{x}=0)\vee(\lambda_{i}^{x}<\mu_{i}^{x})]=0,\forall i\in\{1,2,\cdots,n\}.$
  2. (2)
    For any $(n-1)$-dimensional $L^{0}$-simplex $S_{1}$ contained in $S$ and any $L^{0}$-simplicial subdivision $\mathscr{S}$ of $S_{1}$, $\phi|_{ext_{L^{0}}(\mathscr{S})}$ is a proper $L^{0}$-labeling function of $\mathscr{S}$.
Proof.
Since for any $x\in S$ there exists an unique $\{\lambda_{i}^{x},i=1\sim n\}\subseteq L^{0}(\mathcal{F},[0,1])$ such that $x=\sum_{i=1}^{n}\lambda_{i}^{x}x_{i}$ and $\sum_{i=1}^{n}\lambda_{i}^{x}=1$, where $(\lambda_{1}^{x},\lambda_{2}^{x},\cdots,\lambda_{n}^{x})$ is called the $L^{0}$-barycentric coordinate of $x$. For brevity and clarity, in the proof, we always use $\lambda_{i}^{x}$ for the $i$-th $L^{0}$-barycentric coordinate of $x$ and $\mu_{i}^{x}$ for the $i$-th $L^{0}$-barycentric coordinate of $f(x)$. Then, for any $x\in S$, $x=\sum_{i=1}^{n}\lambda_{i}^{x}x_{i}$ and $f(x)=\sum_{i=1}^{n}\mu_{i}^{x}x_{i}$, we claim that
$\bigvee_{i=1}^{n}[(\lambda_{i}^{x}>0)\wedge(\lambda_{i}^{x}\geq\mu_{i}^{x})]=1.$
Otherwise,
$\bigwedge_{i=1}^{n}[(\lambda_{i}^{x}=0)\vee(\lambda_{i}^{x}<\mu_{i}^{x})]>0.$
Since $\bigvee_{i=1}^{n}(\lambda_{i}^{x}>0)=1$, there exists $j\in\{1,2,\cdots,n\}$ such that
$(\lambda_{j}^{x}>0)\wedge\{\bigwedge_{i=1}^{n}[(\lambda_{i}^{x}=0)\vee(\lambda_{i}^{x}<\mu_{i}^{x})]\}>0,$
namely
$a:=[(\lambda_{j}^{x}>0)\wedge(\lambda_{j}^{x}<\mu_{j}^{x})]\wedge\{\bigwedge_{i\in\{1,2,\cdots,n\}\backslash\{j\}}[(\lambda_{i}^{x}=0)\vee(\lambda_{i}^{x}<\mu_{i}^{x})]\}>0.$
It follows that $I_{a}\lambda_{j}^{x}<I_{a}\mu_{j}^{x}$ and
$I_{a}\lambda_{i}^{x}\leq I_{a}\mu_{i}^{x},\forall i\in\{1,2,\cdots,n\}\backslash\{j\},$
which contradicts with $\sum_{i=1}^{n}\lambda_{i}^{x}=\sum_{i=1}^{n}\mu_{i}^{x}=1$.
For any $x\in S$, let
$b_{i}^{x}=(\lambda_{i}^{x}>0)\wedge(\lambda_{i}^{x}\geq\mu_{i}^{x}),\forall i\in\{1,2,\cdots,n\},$
and further let $a_{1}^{x}=b_{1}^{x}$,
$a_{i}^{x}=b_{i}^{x}\wedge(\bigvee_{j=1}^{i-1}b_{j}^{x})^{c},\forall i\in\{2,3,\cdots,n\}$
and $\phi(x)=\sum_{i=1}^{n}I_{a_{i}^{x}}i$. Since each $a_{i}^{x}$ is uniquely determined by $x$, $f$ and $S$, the mapping $\phi$ defined in this manner is well-defined. Next, we will show that $\phi$ is the desired mapping.
(1). For any $x\in S$, it is clear that $\phi(x)\in L^{0}(\mathcal{F},\{1,2,\cdots,n\})$ and
$(\phi(x)=i)\wedge[(\lambda_{i}^{x}=0)\vee(\lambda_{i}^{x}<\mu_{i}^{x})]=a_{i}^{x}\wedge[(\lambda_{i}^{x}=0)\vee(\lambda_{i}^{x}<\mu_{i}^{x})]=0$
for any $i\in\{1,2,\cdots,n\}$.
(2). For any $\{c_{k},k\in\mathbb{N}\}\in p(1)$ and any sequence $\{y_{k},k\in\mathbb{N}\}$ in $ext_{L^{0}}(\mathscr{S})$, since $y_{k}=\sum_{i=1}^{n}\lambda_{i}^{y_{k}}x_{i}$ and $f(y_{k})=\sum_{i=1}^{n}\mu_{i}^{y_{k}}x_{i}$ for each $k\in\mathbb{N}$, we have
$y:=\sum_{k=1}^{\infty}I_{c_{k}}y_{k}=\sum_{k=1}^{\infty}I_{c_{k}}(\sum_{i=1}^{n}\lambda_{i}^{y_{k}}x_{i})=\sum_{i=1}^{n}(\sum_{k=1}^{\infty}I_{c_{k}}\lambda_{i}^{y_{k}})x_{i}$
and
$f(y):=\sum_{k=1}^{\infty}I_{c_{k}}f(y_{k})=\sum_{k=1}^{\infty}I_{c_{k}}(\sum_{i=1}^{n}\mu_{i}^{y_{k}}x_{i})=\sum_{i=1}^{n}(\sum_{k=1}^{\infty}I_{c_{k}}\mu_{i}^{y_{k}})x_{i},$
then $\lambda_{i}^{y}=\sum_{k=1}^{\infty}I_{c_{k}}\lambda_{i}^{y_{k}}$, $\mu_{i}^{y}=\sum_{k=1}^{\infty}I_{c_{k}}\mu_{i}^{y_{k}}$ and
$b_{i}^{y}$ $=(\lambda_{i}^{y}>0)\wedge(\lambda_{i}^{y}\geq\mu_{i}^{y})$
$=[\bigvee_{k=1}^{\infty}(c_{k}\wedge(\lambda_{i}^{y_{k}}>0))]\wedge[\bigvee_{k=1}^{\infty}(c_{k}\wedge(\lambda_{i}^{y_{k}}\geq\mu_{i}^{y_{k}}))]$
$=\bigvee_{k=1}^{\infty}[c_{k}\wedge(\lambda_{i}^{y_{k}}>0)\wedge(\lambda_{i}^{y_{k}}\geq\mu_{i}^{y_{k}})]$
$=\bigvee_{k=1}^{\infty}(c_{k}\wedge b_{i}^{y_{k}})$
for any $i\in\{1,2,\cdots,n\}$. It follows that
$a_{i}^{y}$ $=b_{i}^{y}\wedge(\bigvee_{j=1}^{i-1}b_{j}^{y})^{c}$
$=[\bigvee_{k=1}^{\infty}(c_{k}\wedge b_{i}^{y_{k}})]\wedge\{\bigvee_{j=1}^{i-1}[\bigvee_{l=1}^{\infty}(c_{l}\wedge b_{j}^{y_{l}})]\}^{c}$
$=\bigvee_{k=1}^{\infty}\{\bigwedge_{j=1}^{i-1}[\bigwedge_{l=1}^{\infty}(c_{k}\wedge b_{i}^{y_{k}})\wedge(c_{l}\wedge b_{j}^{y_{l}})^{c}]\}$
$=\bigvee_{k=1}^{\infty}\{\bigwedge_{j=1}^{i-1}[(c_{k}\wedge b_{i}^{y_{k}})\wedge(c_{k}\wedge b_{j}^{y_{k}})^{c}\wedge(\bigwedge_{l\in\mathbb{N},l\neq k}(c_{k}\wedge b_{i}^{y_{k}})\wedge(c_{l}\wedge b_{j}^{y_{l}})^{c})]\}$
$=\bigvee_{k=1}^{\infty}\{\bigwedge_{j=1}^{i-1}[(c_{k}\wedge b_{i}^{y_{k}})\wedge(b_{j}^{y_{k}})^{c}\wedge(c_{k}\wedge b_{i}^{y_{k}})]\}$
$=\bigvee_{k=1}^{\infty}[c_{k}\wedge b_{i}^{y_{k}}\wedge(\bigvee_{j=1}^{i-1}b_{j}^{y_{k}})^{c}]$
$=\bigvee_{k=1}^{\infty}(c_{k}\wedge a_{i}^{y_{k}})$
for any $i\in\{1,2,\cdots,n\}$. Then, we have
$\phi(\sum_{k=1}^{\infty}I_{c_{k}}y_{k})$ $=\phi(y)$
$=\sum_{i=1}^{n}I_{a_{i}^{y}}i$
$=\sum_{i=1}^{n}I_{\bigvee_{k=1}^{\infty}(c_{k}\wedge a_{i}^{y_{k}})}i$
$=\sum_{i=1}^{n}\sum_{k=1}^{\infty}I_{c_{k}}I_{a_{i}^{y_{k}}}i$
$=\sum_{k=1}^{\infty}I_{c_{k}}(\sum_{i=1}^{n}I_{a_{i}^{y_{k}}}i)$
$=\sum_{k=1}^{\infty}I_{c_{k}}\phi(y_{k}),$
which implies that $\phi$ is $\sigma$-stable, namely $\phi$ is an $L^{0}$-labeling mapping of $\mathscr{S}$.
Moreover, for any $x\in ext_{L^{0}}(\mathscr{S})$, we have
$(\phi(x)=i)\wedge(\lambda_{i}^{x}=0)$ $=a_{i}^{x}\wedge(\lambda_{i}^{x}=0)$
$\leq b_{i}^{x}\wedge(\lambda_{i}^{x}=0)$
$=(\lambda_{i}^{x}>0)\wedge(\lambda_{i}^{x}\geq\mu_{i}^{x})\wedge(\lambda_{i}^{x}=0)$
$=0$
for any $i\in\{1,2,\cdots,n\}$. Thus, $\phi$ is proper. ∎
Lemma 4.9 below is the $L^{0}$-simplex version of random Brouwer fixed point theorem in $RN$ modules.
Lemma 4.9.
Let $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ be an $(n-1)$-dimensional $L^{0}$-simplex in an $RN$ module $E$ and $f:S\rightarrow S$ be a $\sigma$-stable and random sequentially continuous mapping. Then $f$ has a fixed point.
Proof.
Let $\phi$ be the mapping obtained as in Lemma 4.8.
Let $\mathscr{S}$ be an $L^{0}$-barycentric subdivision of $S$. For $\phi$ and $\mathscr{S}$, by Theorem 2.9 there exists an $(n-1)$-dimensional $L^{0}$-simplex $S^{1}:=Conv_{L^{0}}(\{y_{1}^{1},y_{2}^{1},\cdots,y_{n}^{1}\})$, which is a finite concatenation of some finite family of $L^{0}$-simplexes in $\mathscr{S}$, namely there exist a finite family $\{S_{k}^{1},k=1\sim M\}$ of $(n-1)$-dimensional $L^{0}$-simplexes in $\mathscr{S}$ and $\{a_{k}^{1},k=1\sim M\}\in p(1)$ such that
$S^{1}=\sum_{k=1}^{M}I_{a_{k}^{1}}S_{k}^{1}$
is completely labeled by $\phi$ and satisfies that
$\phi(y_{j}^{1})=j~{}\text{for~{}any}~{}j\in\{1,2,\cdots,n\}.$
It is easy to see that
$diam(S^{1})$ $\leq\bigvee_{k=1}^{M}diam(S_{k}^{1})$
$\leq diam(\mathscr{S})$
$\leq\frac{n-1}{n}diam(S).$
Let $\mathscr{S}_{1}$ be an $L^{0}$-barycentric subdivision of $S^{1}$. Now, similarly, for $\phi$ and $\mathscr{S}_{1}$, by Theorem 2.9 there exists an $(n-1)$-dimensional $L^{0}$-simplex $S^{2}:=Conv_{L^{0}}(\{y_{1}^{2},y_{2}^{2},\cdots,y_{n}^{2}\})$, which is a finite concatenation of some finite family of $L^{0}$-simplexes in $\mathscr{S}_{1}$, such that $S^{2}$ is completely labeled by $\phi$ and satisfies that
$\phi(y_{j}^{2})=j~{}\text{for~{}any}~{}j\in\{1,2,\cdots,n\},$
and
$diam(S^{2})\leq\frac{n-1}{n}diam(S^{1})\leq(\frac{n-1}{n})^{2}diam(S).$
By induction, for each $m\in\mathbb{N}$, there exists an $(n-1)$-dimensional $L^{0}$-simplex $S^{m}:=Conv_{L^{0}}(\{y_{1}^{m},y_{2}^{m},\cdots,y_{n}^{m}\})$, which is a finite concatenation of some finite family of $L^{0}$-simplexes in $\mathscr{S}_{m-1}$, such that $S^{m}$ is completely labeled by $\phi$ and satisfies that
$\phi(y_{j}^{m})=j~{}\text{for~{}any}~{}j\in\{1,2,\cdots,n\},$
and
$diam(S^{m})\leq(\frac{n-1}{n})^{m}diam(S),$
where $\mathscr{S}_{m}$ is the $L^{0}$-barycentric subdivision of $S^{m}$ for each $m\in\mathbb{N}$.
Then $\{(y_{1}^{m},y_{2}^{m}\cdots,y_{n}^{m}),m\in\mathbb{N}\}$ is a sequence in $S^{(n)}$, where $S^{(n)}$ is the $n$-th Cartesian power of $S$.
For $\{y_{1}^{m},m\in\mathbb{N}\}$, since $S$ is random sequentially compact by Remark 4.4, there exists a random subsequence $\{y_{1}^{m_{k}^{\prime}},k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to some $y$ in $S$. Since $\{diam(S^{m}),m\in\mathbb{N}\}$ converges $a.s.$ (of course, also in $\mathcal{T}_{\varepsilon,\lambda}$) to 00, by Lemma 3.5 of [26] $\{d^{m_{k}^{\prime}},k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to 00, where
$d^{m_{k}^{\prime}}:=\sum_{l=1}^{\infty}I_{(m_{k}^{\prime}=l)}diam(S^{l}),\forall k\in\mathbb{N}.$
For any $i\in\{2,3,\cdots,n\}$, we have
$\|y_{i}^{m_{k}^{\prime}}-y\|$ $\leq\|y_{i}^{m_{k}^{\prime}}-y_{1}^{m_{k}^{\prime}}\|+\|y_{1}^{m_{k}^{\prime}}-y\|$
$=\sum_{l=1}^{\infty}I_{(m_{k}^{\prime}=l)}\|y_{i}^{l}-y_{1}^{l}\|+\|y_{1}^{m_{k}^{\prime}}-y\|$
$\leq\sum_{l=1}^{\infty}I_{(m_{k}^{\prime}=l)}diam(S^{l})+\|y_{1}^{m_{k}^{\prime}}-y\|$
$=d^{m_{k}^{\prime}}+\|y_{1}^{m_{k}^{\prime}}-y\|,$
which implies that $\{y_{i}^{m_{k}^{\prime}},k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $y$. To sum up, $\{(y_{1}^{m_{k}^{\prime}},y_{2}^{m_{k}^{\prime}},\cdots,y_{n}^{m_{k}^{\prime}})$, $k\in\mathbb{N}\}$ is a random subsequence of $\{(y_{1}^{m},y_{2}^{m},\cdots,y_{n}^{m}),m\in\mathbb{N}\}$ such that $\{y_{i}^{m_{k}^{\prime}},k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $y$ for any $i\in\{1,2,\cdots,n\}$. Since $f$ is random sequentially continuous, by Lemma 4.4 of [26] there exists a random subsequence $\{(y_{1}^{m_{k}},y_{2}^{m_{k}},\cdots,y_{n}^{m_{k}}),k\in\mathbb{N}\}$ of $\{(y_{1}^{m_{k}^{\prime}},y_{2}^{m_{k}^{\prime}},\cdots,y_{n}^{m_{k}^{\prime}}),k\in\mathbb{N}\}$ such that $\{y_{i}^{m_{k}},k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $y$ and $\{f(y_{i}^{m_{k}}),k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $f(y)$ for each $i\in\{1,2,\cdots,n\}$.
Let $y=\sum_{i=1}^{n}\lambda_{i}x_{i}$ and $f(y)=\sum_{i=1}^{n}\mu_{i}x_{i}$, where $\lambda_{i},\mu_{i}\in L_{+}^{0}(\mathcal{F})$ for any $i\in\{1,2,\cdots,n\}$ are such that $\sum_{i=1}^{n}\lambda_{i}=\sum_{i=1}^{n}\mu_{i}=1$. Next, we will show that $f(y)=y$, namely $y$ is a fixed point of $f$.
For any fixed $j\in\{1,2,\cdots,n\}$ and $k\in\mathbb{N}$, let $y_{j}^{m_{k}}=\sum_{i=1}^{n}\lambda_{i}^{m_{k},j}x_{i}$ and $f(y_{j}^{m_{k}})=\sum_{i=1}^{n}\mu_{i}^{m_{k},j}x_{i}$, where $\lambda_{i}^{m_{k},j},\mu_{i}^{m_{k},j}\in L_{+}^{0}(\mathcal{F})$ for any $i\in\{1,2,\cdots,n\}$ and $\sum_{i=1}^{n}\lambda_{i}^{m_{k},j}=\sum_{i=1}^{n}\mu_{i}^{m_{k},j}=1$. Since $\phi$ is $\sigma$-stable, we have
$\phi(y_{j}^{m_{k}})$ $=\phi(\sum_{l=1}^{\infty}I_{(m_{k}=l)}y_{j}^{m^{\prime}_{l}})$
$=\sum_{l=1}^{\infty}I_{(m_{k}=l)}\phi(y_{j}^{m^{\prime}_{l}})$
$=\sum_{l=1}^{\infty}I_{(m_{k}=l)}\phi(\sum_{r=1}^{\infty}I_{(m^{\prime}_{l}=r)}y_{j}^{r}))$
$=\sum_{l=1}^{\infty}I_{(m_{k}=l)}\sum_{r=1}^{\infty}I_{(m^{\prime}_{l}=r)}\phi(y_{j}^{r})$
$=\sum_{l=1}^{\infty}I_{(m_{k}=l)}\sum_{r=1}^{\infty}I_{(m^{\prime}_{l}=r)}j$
$=j,$
then $\lambda_{j}^{m_{k},j}\geq\mu_{j}^{m_{k},j}$ by Lemma 4.8.
Furthermore, by Lemma 5.5 of [26], $\{\lambda_{j}^{m_{k},j},k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $\lambda_{j}$ and $\{\mu_{j}^{m_{k},j},k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $\mu_{j}$ for each $j\in\{1,2,\cdots,n\}$, then
$\lambda_{j}\geq\mu_{j},~{}\forall j\in\{1,2,\cdots,n\}.$
However, $\sum_{j=1}^{n}\lambda_{j}=\sum_{j=1}^{n}\mu_{j}=1$, then
$\lambda_{j}=\mu_{j},~{}\forall j\in\{1,2,\cdots,n\}.$
Thus, $f(y)=y$. ∎
Remark 4.10.
Lemma 4.9 is a generalization of [11, Theorem 2.3], since $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is an $RN$ module, and a $\sigma$-stable a.s. sequentially continuous mapping is, obviously, random sequentially continuous, as noted in Remark 4.6. Since random sequential continuity is much more complicated than a.s. sequential continuity, Lemma 4.4 of [26] has played a key role in the treatment of random sequential continuity in the proof of Lemma 4.9. Besides, we would also like to emphasize another difference between the proofs of Lemma 4.9 and [11, Theorem 2.3]: the sequence $\{S^{m},m\in\mathbb{N}\}$ of $(n-1)$-dimensional $L^{0}$-simplexes, which is constructed in Lemma 4.9, has the nice property that $S^{m+1}\subseteq S^{m}$ for each $m\in\mathbb{N}$, but the sequence $\{S^{m},m\in\mathbb{N}\}$ constructed in [11, Theorem 2.3] does not have this property. The nice property of our sequence will be used to study other problems in the future. Besides, the proof of Lemma 4.9 is simpler and clearer than that of [11, Theorem 2.3] since we only employ a single proper $L^{0}$-labeling function throughout the proof of Lemma 4.9.
Let $H$ be a Hilbert space and $G$ be a nonempty closed convex subset of $H$, then it is well known that the classical projection $P_{G}$ of $H$ onto $G$ is nonexpensive. By a completely similar method, one can prove the following lemma:
Lemma 4.11.
Let $(H,\langle\cdot,\cdot\rangle)$ be a $\mathcal{T}_{\varepsilon,\lambda}$-complete random inner product module over $\mathbb{K}$ with base $(\Omega,\mathcal{F},P)$ and $G$ be a nonempty $\mathcal{T}_{\varepsilon,\lambda}$-closed $L^{0}$-convex subset of $H$. Then, for any $x\in H$, there exists an unique element $P_{G}(x)$ in $G$ such that
$\|x-P_{G}(x)\|=\bigwedge\{\|x-g\|:g\in G\},$
[…]
$\|P_{G}(x)-P_{G}(y)\|\leq\|x-y\|,\forall x,y\in H.$
In Lemma 4.11, since $G$ is $\sigma$-stable and $P_{G}$ is $L^{0}$-Lipschitzian, then $P_{G}$ is $\sigma$-stable by Lemma 2.11 of [27].
We can now prove Theorem 2.1.
Proof of Theorem 2.1.
Since $G$ is $a.s.$ bounded, there exists an $L^{0}$-simplex $S$ in $L^{0}(\mathcal{F},\mathbb{R}^{d})$ such that $G\subseteq S$. Since $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is a $\mathcal{T}_{\varepsilon,\lambda}$-complete random inner product module, let $h:S\rightarrow G$ be the restriction of $P_{G}$ to $S$, where $P_{G}$ is the random projection of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ onto $G$, it is clear that $h$ is $\mathcal{T}_{c}$-continuous.
Consider the mapping $\hat{f}:S\rightarrow G\subseteq S$ defined by
$\hat{f}(x)=(f\circ h)(x),\forall x\in S.$
Since $f$ is $\sigma$-stable and $\mathcal{T}_{c}$-continuous, it follows that $\hat{f}$ is also $\sigma$-stable and $\mathcal{T}_{c}$-continuous. By Remark 4.6, $\hat{f}$ is $\sigma$-stable and random sequentially continuous, then, by Lemma 4.9 there exists $z\in S$ such that $\hat{f}(z)=z$, which implies that $z\in G$. By the definition of $h$, we have $h(z)=z$, and hence $f(z)=z$. ∎
Proposition 4.12 below was first established as Lemma 4.7 of [26], which is merely a reformulation of Theorem 2.1 in the case of a finitely generated $RN$ module. For the proof of Proposition 4.12, we refer the reader to [26].
Proposition 4.12.
Let $G$ be an $L^{0}$-convex, $\mathcal{T}_{\varepsilon,\lambda}$-closed and $a.s.$ bounded subset of a finitely generated $RN$ module $(E,\|\cdot\|)$ over $\mathbb{K}$ with base $(\Omega,\mathcal{F},P)$ and $f:G\rightarrow G$ a $\sigma$-stable and random sequentially continuous mapping, where we recall that $E$ is finitely generated if there exists a finite subset of $E$, denoted by $\{x_{1},x_{2},\cdots,x_{d}\}$ for some $d\in\mathbb{N}$, such that $E=\{\sum^{d}_{i=1}\xi_{i}\cdot x_{i}:\xi_{i}\in L^{0}(\mathcal{F},\mathbb{K})$ for each $i\in\{1,2,\cdots,d\}\}$. Then $f$ has a fixed point.
5 The equivalence of Proposition 2.3 and Proposition 2.4 together with other interesting results
Lemma 5.1.
Let $G$ be a $\sigma$-stable subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$. Then a mapping $h:G\rightarrow G$ is local iff $h$ is $\sigma$-stable.
Proof.
Necessity. For any sequence $\{x_{n},n\in\mathbb{N}\}$ in $G$ and any countable partition $\{a_{n},n\in\mathbb{N}\}$ of unity in $B_{\mathcal{F}}$, then, according to the local property of $h$, $I_{a_{n}}h(\sum_{n=1}^{\infty}I_{a_{n}}x_{n})=I_{a_{n}}h(x_{n})$, so that
$h(\sum_{n=1}^{\infty}I_{a_{n}}x_{n})=(\sum_{n=1}^{\infty}I_{a_{n}})h(\sum_{n=1}^{\infty}I_{a_{n}}x_{n})=\sum_{n=1}^{\infty}I_{a_{n}}h(x_{n}).$
Sufficiency. For any $a\in B_{\mathcal{F}}$ and any $x,y$ in $G$, from the $\sigma$-stability of $h$ one has that $h(I_{a}x+I_{a^{c}}y)=I_{a}h(x)+I_{a^{c}}h(y)$. Then, when $I_{a}x=I_{a}y$, $h(y)=I_{a}h(x)+I_{a^{c}}h(y)$, so $I_{a}h(x)=I_{a}h(y)$, which amounts to saying that $h$ is local. ∎
Lemma 5.2 below is a special case of the main result of [42], which is enough for our purpose.
Lemma 5.2 ([42]).
Let $X$ and $Y$ be two complete separable metric spaces, $(\Omega,\mathcal{F},P)$ a complete probability space and $h:L^{0}(\mathcal{F},X)\rightarrow L^{0}(\mathcal{F},Y)$ a local and continuous-in-probability operator. Then there exists a Carathéodory function $f:\Omega\times X\rightarrow Y$ (namely $f(\cdot,x)$ is measurable for any $x\in X$ and $f(\omega,\cdot)$ is continuous for any $\omega\in\Omega$) such that $h=h_{f}$. Here, $h_{f}:L^{0}(\mathcal{F},X)\rightarrow L^{0}(\mathcal{F},Y)$ is defined by $h_{f}(x)=$ the equivalence class of $f(\cdot,x^{0}(\cdot))$, where $x^{0}$ is an arbitrarily chosen representative of $x$.
In fact, Lemma 5.2 has implied that $h$ is a.s. sequentially continuous, namely $\{h(x_{n}),n\in\mathbb{N}\}$ converges a.s. to $h(x)$ whenever $\{x_{n},n\in\mathbb{N}\}$ converges a.s. to $x$.
Lemma 5.3.
Let $U:\Omega\rightarrow 2^{\mathbb{R}^{d}}$ be the same as in Proposition 2.4, then $G_{U}:=\{x\in L^{0}(\mathcal{F},\mathbb{R}^{d}):x$ has a representative $x^{0}$ such that $x^{0}(\omega)\in U(\omega)~{}a.s.\}$ is an a.s. bounded $\mathcal{T}_{\varepsilon,\lambda}$-closed $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$.
Proof.
[…]
Method 1. First, let us define a real-valued function $r:\Omega\rightarrow[0,+\infty)$ by $r(\omega)=\sup\{|x|:x\in U(\omega)\}$ for any $\omega\in\Omega$, then $r$ is well-defined since each $U(\omega)$ is bounded. Now, let $G_{U}^{0}=\{x^{0}:x^{0}$ is an $\mathbb{R}^{d}$-valued $\mathcal{F}$-measurable function on $\Omega$ such that $x^{0}(\omega)\in U(\omega)~{}a.s.\}$, then $|x^{0}(\omega)|\leq r(\omega)$ a.s. for each $x^{0}\in G_{U}^{0}$. Further, let $\xi^{0}=ess.sup\{|x^{0}|:x^{0}\in G_{U}^{0}\}$, see, for example, [23, 26] for the notion of an essential supremum for a set of extended real-valued random variables, then it still holds that $\xi^{0}(\omega)\leq r(\omega)$ a.s., so we can assume that $\xi^{0}$ is a real-valued random variable. Then $\bigvee\{\|x\|:x\in G_{U}\}\leq\xi$, where $\xi$ is the equivalence class of $\xi^{0}$, namely $G_{U}$ is a.s. bounded.
Method 2. Let $\hat{\mathcal{F}}$ be the completion of $\mathcal{F}$ with respect to $P$, then $Gr(U)\in\hat{\mathcal{F}}\otimes Bor(\mathbb{R}^{d})$ and hence by Theorem 4.2 of [51] there exists a sequence $\{x_{n}^{0},n\in\mathbb{N}\}$ of $\hat{\mathcal{F}}$-measurable selections of $U$ such that $U(\omega)=\overline{\{x_{n}^{0}(\omega):n\in\mathbb{N}\}}$ for each $\omega\in\Omega$. Let $r^{0}:\Omega\rightarrow[0,+\infty)$ be the function defined by $r^{0}(\omega)=\sup\{|x_{n}^{0}(\omega)|:n\in\mathbb{N}\}$ for any $\omega\in\Omega$, then $r^{0}$ is $\hat{\mathcal{F}}$-measurable and hence there exists an $\mathcal{F}$-measurable nonnegative random variable $r_{1}^{0}$ such that $r^{0}(\omega)=r_{1}^{0}(\omega)$ a.s. Further, let $r_{1}$ be the equivalence class of $r_{1}^{0}$, then it is clear that $\bigvee\{\|x\|:x\in G_{U}\}\leq r_{1}$, $G_{U}$ is again a.s. bounded. ∎
Theorem 5.4.
Proposition 2.3 is equivalent to Proposition 2.4.
Proof.
Proposition 2.3 $\Rightarrow$ Proposition 2.4 is clear by Lemma 5.1 and the fact that the $\mathcal{T}_{\varepsilon,\lambda}$-continuity of $h$ is exactly the continuity-in-probability of $h$.
For the reverse implication, for any $\xi\in L_{++}^{0}(\mathcal{F})$, let $T_{\xi}$ be the isomorphism of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ to itself, defined by $T_{\xi}(x)=\xi x$ for any $x\in L^{0}(\mathcal{F},\mathbb{R}^{d})$. Then, it is also clear that $T_{\xi}^{-1}=T_{\xi^{-1}}$. Now, let $G$ and $f$ be the same as assumed in Proposition 2.3, then there exists some $\xi\in L_{++}^{0}(\mathcal{F})$ such that $\xi^{-1}G\subseteq L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, where $\bar{B}_{1}(\mathbb{R}^{d})=\{y\in\mathbb{R}^{d}:|y|\leq 1\}$ and ||$\cdot$|| ⋅ | stands for the Euclidean norm on $\mathbb{R}^{d}$. For brevity, we denote $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ by $V$, then $V$ is an a.s. bounded, $\mathcal{T}_{\varepsilon,\lambda}$-closed and $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$. Further, let $P_{V,\xi^{-1}G}$ be the restriction of $P_{\xi^{-1}G}$ to $V$, where $P_{\xi^{-1}G}$ is the random projection of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ onto $\xi^{-1}G$, then $h:=T_{\xi^{-1}}\circ f\circ T_{\xi}\circ P_{V,\xi^{-1}G}$ is a $\sigma$-stable and $\mathcal{T}_{\varepsilon,\lambda}$-continuous mapping from $V$ to $V$, and hence $h$ is also a local and continuous-in-probability operator. Now, define a set-valued function $U:\Omega\rightarrow 2^{\mathbb{R}^{d}}$ by $U(\omega)=\bar{B}_{1}(\mathbb{R}^{d})$ for each $\omega\in\Omega$, then $U$ not only satisfies the hypothesis of Proposition 2.4 but also is weakly measurable in the sense of [51], and in particular $G_{U}$ is exactly $V$. Thus, there exists $v\in V$ such that $h(v)=v$ by Proposition 2.4, then $\xi v=f(\xi P_{V,\xi^{-1}G}(v))\in G$. Again, let $g\in G$ be such that $g=\xi v$, then $v=\xi^{-1}g\in\xi^{-1}G$, then $P_{V,\xi^{-1}G}(v)=\xi^{-1}g$, and hence $f(\xi P_{V,\xi^{-1}G}(v))=f(g)=\xi v=g$, namely $g$ is a fixed point of $f$ in $G$. ∎
Remark 5.5.
Let $V=L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ and $h:V\rightarrow V$ be the same as obtained in the process of the proof of Proposition 2.4 implying Proposition 2.3. If $(\Omega,\mathcal{F},P)$ is a complete probability space, then by Lemma 5.2 there exists a Carathéodory function $f:\Omega\times\bar{B}_{1}(\mathbb{R}^{d})\rightarrow\bar{B}_{1}(\mathbb{R}^{d})$ such that $h(v)=h_{f}(v)$ for any $v\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$. This observation leads us to the following interesting (or even somewhat surprising) result.
Theorem 5.6.
Both Proposition 2.3 and Proposition 2.4 are equivalent to either of the following two assertions:
  1. (1)
    Every $\mathcal{T}_{\varepsilon,\lambda}$-continuous $\sigma$-stable mapping $h$ from $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ to itself has a fixed point.
  2. (2)
    Every Carathéodory function function $f:(\Omega,\mathcal{F},P)\times\bar{B}_{1}(\mathbb{R}^{d})\rightarrow\bar{B}_{1}(\mathbb{R}^{d})$ admits a random fixed point, namely there exists a random vector $v:(\Omega,\mathcal{F},P)\rightarrow\bar{B}_{1}(\mathbb{R}^{d})$ such that $f(\omega,v(\omega))=v(\omega)$ a.s..
Proof.
Theorem 5.4 has shown that Proposition 2.3 and Proposition 2.4 are equivalent, while the second part of the proof of Theorem 5.4 has also shown that Proposition 2.3 is equivalent to (1). Thus, we only need to prove that (1) and (2) are equivalent.
$(1)\Rightarrow(2)$. Let $h=h_{f}$, then there exists some $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ such that $h_{f}(x)=x$, and hence taking $v$ to be an arbitrarily chosen representative of $x$ yields $f(\omega,v(\omega))=v(\omega)$ a.s., namely $v$ satisfies (2).
$(2)\Rightarrow(1)$. Let $\hat{\mathcal{F}}$ be the completion of $\mathcal{F}$ with respect to $P$ and further notice the fact that $L^{0}(\hat{\mathcal{F}},\bar{B}_{1}(\mathbb{R}^{d}))=L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ although $\hat{\mathcal{F}}$ and $\mathcal{F}$ may differ, then by Lemma 5.2 there exists some Carathéodory function $f:(\Omega,\hat{\mathcal{F}},P)\times\bar{B}_{1}(\mathbb{R}^{d})\rightarrow\bar{B}_{1}(\mathbb{R}^{d})$ such that $h=h_{f}$. By (1) there exists some $\hat{\mathcal{F}}$-measurable $v:\Omega\rightarrow\bar{B}_{1}(\mathbb{R}^{d})$ such that $f(\omega,v(\omega))=v(\omega)$ a.s.. Let $x$ be the equivalence class of $v$, then $h(x)=h_{f}(x)=x$, since $L^{0}(\hat{\mathcal{F}},\bar{B}_{1}(\mathbb{R}^{d}))=L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, $x$ is also regarded as a fixed point of $h$ in $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$. ∎
Remark 5.7.
The original Brouwer fixed point theorem states that every continuous function $f$ from $\bar{B}_{1}(\mathbb{R}^{d})$ to itself has a fixed point, see [3]. (2) of Theorem 5.6 can be aptly called the random version of the original Brouwer fixed point, which is equivalent to either of Proposition 2.3 and Proposition 2.4. Thus, why we can obtain such a deep result as Theorem 5.6 may lie in the fact that Lemma 5.2 due to Ponosov is surprisingly deep! Proposition 5.8 below should also be credited to Lemma 5.2.
Proposition 5.8.
Let $G$ be a $\mathcal{T}_{\varepsilon,\lambda}$-closed $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ and $T:G\rightarrow L^{0}(\mathcal{F},\mathbb{R}^{d})$ be a $\sigma$-stable $\mathcal{T}_{\varepsilon,\lambda}$-continuous mapping. Then $T$ must be a.s. sequentially continuous.
Proof.
Since elements of both $G$ and $L^{0}(\mathcal{F},\mathbb{R}^{d})$ are equivalence classes, we can, without loss of generality, assume that $(\Omega,\mathcal{F},P)$ is complete. Further, let $P_{G}$ be the random projection of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ onto $G$, then $T\circ P_{G}:L^{0}(\mathcal{F},\mathbb{R}^{d})\rightarrow L^{0}(\mathcal{F},\mathbb{R}^{d})$ is still $\sigma$-stable and $\mathcal{T}_{\varepsilon,\lambda}$-continuous, and hence by Lemma 5.2 there exists some Carathéodory function $f:(\Omega,\mathcal{F},P)\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ such that $T\circ P_{G}=h_{f}$, in particular $T(g)=h_{f}(g)$ for any $g\in G$, which has implied that $T$ is a.s. sequentially continuous. ∎
Remark 5.9.
Proposition 5.8 has implied the equivalence of Proposition 3.1 of [11] with either of Proposition 2.3 and Proposition 2.4.
6 The equivalence of random Brouwer fixed point theorem and random Borsuk theorem
Proposition 6.1 below shows that Theorem 2.1 is equivalent to a special case of it, namely random Brouwer fixed point theorem for random closed balls, which will provide much convenience for our research into random Borsuk theorem.
Proposition 6.1.
The following three statements are equivalent:
  • (1)
    Every $\mathcal{T}_{c}$-continuous $\sigma$-stable mapping from a $\mathcal{T}_{c}$-closed, $\sigma$-stable, a.s. bounded and $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ to itself has a fixed point.
  • (2)
    Every $\mathcal{T}_{c}$-continuous $\sigma$-stable mapping from a random closed ball $\bar{B}(\theta,r)$ of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ to itself has a fixed point, where $r\in L^{0}_{++}(\mathcal{F})$ and $\bar{B}(\theta,r)=\{x\in L^{0}(\mathcal{F},\mathbb{R}^{d}):\|x\|\leq r\}$.
  • (3)
    Every $\mathcal{T}_{c}$-continuous $\sigma$-stable mapping from the random closed unit ball $\bar{B}(\theta,1)$ of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ to itself has a fixed point.
Proof.
$(1)\Rightarrow(2)$ is clear.
$(2)\Rightarrow(1)$. Let $G$ be a $\mathcal{T}_{c}$-closed, $\sigma$-stable, a.s. bounded and $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ and $f:G\rightarrow G$ be a $\mathcal{T}_{c}$-continuous $\sigma$-stable mapping. Since $G$ is a.s. bounded, we can choose an enough large $r\in L^{0}_{++}(\mathcal{F})$ such that $G\subseteq\bar{B}(\theta,r)$. Further, since $G$ is $\sigma$-stable and $\mathcal{T}_{c}$-closed, it follows immediately from [23] that $G$ is also $\mathcal{T}_{\varepsilon,\lambda}$-closed, then by Lemma 4.11 the random projection $P_{G}$ of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ onto $G$ is well defined, $\sigma$-stable and $\mathcal{T}_{c}$-continuous. Now, let $P_{\bar{B}(\theta,r),G}$ be the restriction of $P_{G}$ to $\bar{B}(\theta,r)$, then it is also $\sigma$-stable and $\mathcal{T}_{c}$-continuous since $\bar{B}(\theta,r)$ is $\sigma$-stable. Thus, it is clear that $f\circ P_{\bar{B}(\theta,r),G}$ is a $\mathcal{T}_{c}$-continuous and $\sigma$-stable mapping from $\bar{B}(\theta,r)$ to $\bar{B}(\theta,r)$, then by (2) $f\circ P_{\bar{B}(\theta,r),G}$ has a fixed point $x$. Eventually, $f(x)=x$ since $x\in G$.
$(2)\Leftrightarrow(3)$ is clear since, $T_{r}:\bar{B}(\theta,1)\rightarrow\bar{B}(\theta,r)$ defined by $T_{r}(x)=rx$ for each $x\in\bar{B}(\theta,1)$, is a $\sigma$-stable $\mathcal{T}_{c}$-homeomorphism. ∎
In 1967, Borsuk [2] introduced the notion of an $r$-image for a subset of $\mathbb{R}^{d}$. Now, according to the two topologies $\mathcal{T}_{c}$ and $\mathcal{T}_{\varepsilon,\lambda}$ of $L^{0}(\mathcal{F},\mathbb{R}^{d})$, we can introduce the corresponding two notions of a $\mathcal{T}_{c}$-$r$-image and a $\mathcal{T}_{\varepsilon,\lambda}$-$r$-image for a $\sigma$-stable subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$, respectively, as follows.
Definition 6.2.
Let $E$ and $F$ be two $\sigma$-stable subsets of $L^{0}(\mathcal{F},\mathbb{R}^{d})$. $E$ is called a $\mathcal{T}_{c}$-$r$-image (resp., a $\mathcal{T}_{\varepsilon,\lambda}$-$r$-image) of $F$ if there exist a $\mathcal{T}_{c}$-continuous (resp., $\mathcal{T}_{\varepsilon,\lambda}$-continuous) $\sigma$-stable mapping $h:F\rightarrow E$ and a $\mathcal{T}_{c}$-continuous (resp., $\mathcal{T}_{\varepsilon,\lambda}$-continuous) $\sigma$-stable mapping $g:E\rightarrow F$ such that $h\circ g$ is the identity on $E$. Such a mapping $h$ is called a $\mathcal{T}_{c}$-$r$-mapping (resp., a $\mathcal{T}_{\varepsilon,\lambda}$-$r$-mapping) of $F$ onto $E$. In the special case where $E\subseteq F$ and $g$ is the inclusion mapping, then $E$ is called a $\mathcal{T}_{c}$-retract (resp., a $\mathcal{T}_{\varepsilon,\lambda}$-retract) of $F$ and $h$ is called a $\mathcal{T}_{c}$-retraction (resp., a $\mathcal{T}_{\varepsilon,\lambda}$-retraction).
Let $E$ and $F$ be the same as in Definition 6.2. If $h:F\rightarrow E$ is a $\sigma$-stable and bijective mapping, then it is easy to check that $h^{-1}:E\rightarrow F$ is also $\sigma$-stable. It is also clear that a $\sigma$-stable $\mathcal{T}_{c}$-homeomorphism (resp., $\mathcal{T}_{\varepsilon,\lambda}$-homeomorphism) of $F$ onto $E$ must be a $\mathcal{T}_{c}$-$r$-mapping (resp., a $\mathcal{T}_{\varepsilon,\lambda}$-$r$-mapping) of $F$ onto $E$. It follows immediately from Lemma 4.3 of [26] that a $\mathcal{T}_{\varepsilon,\lambda}$-$r$-mapping (resp., a $\mathcal{T}_{\varepsilon,\lambda}$-retraction) of $F$ onto $E$ must be a $\mathcal{T}_{c}$-$r$-mapping (resp., a $\mathcal{T}_{c}$-retraction) of $F$ onto $E$.
Proposition 6.3 below is a random version of Theorem 6.9 of [1], which is essentially a reformulation of Theorem 2.1 in a more general form.
Proposition 6.3.
Let $E$ be a a $\mathcal{T}_{c}$-$r$-image of an a.s. bounded, $\mathcal{T}_{c}$-closed, $\sigma$-stable and $L^{0}$-convex subset $F$ of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ and $f:E\rightarrow E$ be $\sigma$-stable and $\mathcal{T}_{c}$-continuous. Then $f$ has a fixed point.
Proof.
Let $g:E\rightarrow F$ and $h:F\rightarrow E$ be the same as in Definition 6.2, then $g\circ f\circ h:F\rightarrow F$ has a fixed point $z$ by Theorem 2.1. Let $x=h(z)$, then
$f(x)=(h\circ g)(f(x))=(h\circ g\circ f\circ h)(z)=h(z)=x.$
Theorem 6.4 below can be aptly called random Borsuk theorem, which is a random version of the classical Borsuk theorem [2] (see also [1, Theorem 6.12]). Here, we would like to remind the reader of the fact that $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ in the sequel of this section is the same thing as the random closed unit ball $\bar{B}(\theta,1):=\{x\in L^{0}(\mathcal{F},\mathbb{R}^{d}):\|x\|\leq 1\}$.
Theorem 6.4.
$L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$ is not a $\mathcal{T}_{c}$-$r$-image of $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, where, as usual, $\bar{B}_{1}(\mathbb{R}^{d})=\{y\in\mathbb{R}^{d}:|y|\leq 1\}$, $\partial\bar{B}_{1}(\mathbb{R}^{d})=\{y\in\mathbb{R}^{d}:|y|=1\}$ and ||$\cdot$|| ⋅ | denotes the Euclidean norm on $\mathbb{R}^{d}$. In particular, $L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$ is not a $\mathcal{T}_{c}$-retraction of $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$.
Proof.
It is easy to see that both $L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$ and $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ are $\sigma$-stable. If $L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$ is a $\mathcal{T}_{c}$-$r$-image of $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, then there are a $\sigma$-stable $\mathcal{T}_{c}$-continuous mapping $h:L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$ and a $\sigma$-stable $\mathcal{T}_{c}$-continuous mapping $g:L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ such that $h\circ g$ is the identity on $L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$. Define $f:L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ by $f(x)=g(-h(x))$ for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, then $f$ is a $\sigma$-stable and $\mathcal{T}_{c}$-continuous mapping, and hence there exists $z\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ by Theorem 2.1 such that $f(z)=z$, that is to say, $z=g(-h(z))$, so $h(z)=(h\circ g)(-h(z))=-h(z)$. Thus $h(z)=0\in L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$, a contradiction. ∎
Corollary 6.5 below is essentially Lemma 3.5 of [11].
Corollary 6.5.
$L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$ is not a $\mathcal{T}_{\varepsilon,\lambda}$-$r$-image of $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$. In particular, $L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$ is not a $\mathcal{T}_{\varepsilon,\lambda}$-retraction of $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$.
Theorem 6.6.
Theorem 6.4 is equivalent to random Brouwer fixed point theorem for the random closed unite ball $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ (and hence also equivalent to Theorem 2.1 by Proposition 6.1).
Proof.
The proof of Theorem 6.4 has shown that the random Brouwer fixed point theorem implies Theorem 6.4. We only need to prove that Theorem 6.4 implies the random Brouwer fixed point theorem for the random closed unit ball of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ as follows.
Suppose that there exists a $\sigma$-stable and $\mathcal{T}_{c}$-continuous mapping $f:L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ such that $f$ has no fixed point. Define $e:L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F})$ by $e(x)=\|f(x)-x\|$ for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, and denote $\eta=\bigwedge\{e(x):x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\}$, then by Lemma 3.4 of [26] there exists $x_{0}\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ such that $\eta=\|f(x_{0})-x_{0}\|$ since $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\}$ is random sequentially compact and $\sigma$-stable and since $e$ is $\sigma$-stable and random sequentially continuous. Then $\eta>0$, since otherwise, $\eta=0$ yields $f(x_{0})=x_{0}$, which contradicts the assumption that $f$ has no fixed point.
Now, let $A\in\mathcal{F}$ be an arbitrary chosen representative of $(\eta>0)$, then $P(A)>0$. Since $\theta\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ and $f:L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ is $\sigma$-stable, then $f$ has the local property, namely $\tilde{I}_{C}f(x)=\tilde{I}_{C}f(\tilde{I}_{C}x)$ for any $C\in\mathcal{F}$ and $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$: in fact, since
$f(\tilde{I}_{C}x)=f(\tilde{I}_{C}x+\tilde{I}_{C^{c}}\theta)=\tilde{I}_{C}f(x)+\tilde{I}_{C^{c}}f(\theta),$
then $\tilde{I}_{C}f(\tilde{I}_{C}x)=\tilde{I}_{C}f(x)$. In particular, $\tilde{I}_{A}f(\tilde{I}_{A}x)=\tilde{I}_{A}f(x)$ for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$. Further, let
$G_{A}=\tilde{I}_{A}L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d})):=\{\tilde{I}_{A}x:x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\}$
and define $f_{A}:G_{A}\rightarrow G_{A}$ by $f_{A}(\tilde{I}_{A}x)=\tilde{I}_{A}f(x)$ for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, then $f_{A}$ is well defined. It is clear that $G_{A}$ is $\sigma$-stable, $\mathcal{T}_{c}$-closed, a.s. bounded and $L^{0}(\mathcal{F})$-convex subset of $L^{0}(\mathcal{F},R^{d})$ and $f_{A}$ is also $\sigma$-stable and $\mathcal{T}_{c}$-continuous.
Let $(A,A\cap\mathcal{F},P_{A})$ be the probability space with $P_{A}:A\cap\mathcal{F}\rightarrow[0,1]$ defined by $P_{A}(A\cap C)=\frac{P(A\cap C)}{P(A)}$ for any $C\in\mathcal{F}$. Further, identifying $G_{A}$ with $L^{0}(\mathcal{F}_{A},\bar{B}_{1}(\mathbb{R}^{d}))$, where $\mathcal{F}_{A}=A\cap\mathcal{F}$, then $f_{A}$ can be also identified with a mapping from $L^{0}(\mathcal{F}_{A},\bar{B}_{1}(\mathbb{R}^{d}))$ to $L^{0}(\mathcal{F}_{A},\bar{B}_{1}(\mathbb{R}^{d}))$. Again observing $\|f(x)-x\|>0$ on $A$ for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, namely $\|f_{A}(x)-x\|>0$ on $A$ for each $x\in G_{A}$, we can, without loss of generality, assume that $A=\Omega$, then $\|f(x)-x\|>0$ on $\Omega$ for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ (since otherwise, we can consider $f_{A}$ and $L^{0}(\mathcal{F}_{A},\bar{B}_{1}(\mathbb{R}^{d}))$ to replace $f$ and $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$).
Finally, let $L_{x}=\{\lambda\in L^{0}(\mathcal{F}):\|x+\lambda(f(x)-x)\|=1\}$ for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ and $\lambda(x)=Min~{}L_{x}:=$ the smallest element of $L_{x}$. In fact, $L_{x}$ has only the two elements, which satisfy the equation
$\|x\|^{2}+2\lambda\langle x,f(x)-x\rangle+\lambda^{2}\|f(x)-x\|^{2}-1=0,$
where $\langle x,y\rangle$ stands for the $L^{0}$-inner product of $x$ and $y$ in $L^{0}(\mathcal{F},\mathbb{R}^{d})$. Similar to the Vieta theorem, one can have
$\lambda(x)=\frac{\langle x,x-f(x)\rangle-\sqrt{\langle x,x-f(x)\rangle^{2}+\|x-f(x)\|^{2}(1-\|x\|)^{2}}}{\|f(x)-x\|^{2}}.$
This define a mapping $\lambda:L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F})$ and it is not very difficult to verify that $\lambda$ is still $\sigma$-stable and $\mathcal{T}_{c}$-continuous. By noticing that $\langle x,x-f(x)\rangle=\|x\|^{2}-\langle x,f(x)\rangle$ and the random Schwartz inequality: $|\langle x,f(x)\rangle|\leq\|x\|\|f(x)\|$, then
$\langle x,x-f(x)\rangle\geq\|x\|(1-\|f(x)\|)\geq 0$
for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$. In particular, when $x\in L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$, namely when $\|x\|=1$, one has $\lambda(x)=0$. Again, define $h:L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$ by $h(x)=x+\lambda(x)(f(x)-x)$, then $h$ is a $\mathcal{T}_{c}$-retraction of $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ onto $L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$, which contradicts Theorem 6.4. ∎
Remark 6.7.
Since $h$ obtained in the final part of the proof of Theorem 6.6 is a $\mathcal{T}_{c}$-retraction. Actually, we have obtained an interesting conclusion, namely the random Brouwer fixed point theorem is equivalent to the fact that $L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$ is not a $\mathcal{T}_{c}$-retraction of $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$!
Remark 6.8.
By a completely similar method used in Theorem 6.6, one can easily prove that Corollary 6.5 is equivalent to the $\mathcal{T}_{\varepsilon,\lambda}$-version of random Brouwer fixed point theorem.
Acknowledgements The authors would like to thank Professor Hong-Kun Xu for his valuable comments on this work.
Funding This work was supported by the National Natural Science Foundation of China (Grant No.12371141) and the Natural Science Foundation of Hunan Province of China (Grant No.2023JJ30642).
Data availability. Data sharing not applicable to this article as no datasets were generated or analyzed during the current study.
Declarations
Conflict of interest There is no conflict of interest.